### v1

# Learning to combine top-down context and feed-forward representations under ambiguity with apical and basal dendrites

Abstract
One of the most striking features of neocortical anatomy is the presence of extensive top-down projections into primary sensory areas. Notably, many of these top-down projections impinge on the distal apical dendrites of pyramidal neurons, where they exert a modulatory effect, altering the gain of responses. It is thought that these top-down projections carry contextual information that can help animals to resolve ambiguities in sensory data. However, it has yet to be demonstrated how such modulatory connections to the distal apical dendrites can serve this computational function. Here, we develop a computational model of pyramidal cells that integrates contextual information from top-down projections to apical compartments with sensory representations driven by bottom-up projections to basal compartments. When input stimuli are ambiguous and relevant contextual information is available, the apical feedback modulates the basal signals to recover unambiguous sensory representations. Importantly, when stimuli are unambiguous, contextual information which is irrelevant or opposes sensory evidence is appropriately ignored by the model. By generalizing the task to temporal sequences, we further show that our model can learn to integrate contextual information across time. Using layer-wise relevance propagation, we extract the importance of individual neurons to the prediction of each category, revealing that neurons that are most relevant for the overlap of categories receive the largest magnitude of top-down signals, and are necessary for solving the task. This work thus provides a proof-of-concept demonstrating how the top-down modulatory inputs to apical dendrites in sensory regions could be used by the cortex to handle the ambiguities that animals encounter in the real world.
1 Introduction
Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information [94, 88]. Contextual information can be derived from a number of sources, including the spatial and temporal domains [65], task demands [69], as well as other sensory modalities.
Several key neuroanatomical and cellular features could support the computations associated with context integration. In the mammalian brain, context has been proposed to be conveyed by top-down feedback pathways which are abundant in sensory regions of the neocortex [66, 85, 72]. Perturbations of top-down pathways from higher order areas have been associated with delayed object recognition [75] as well as disrupted stimulus response curves in lower-order regions [104, 90], which may reflect deficits in contextual processing. Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images [68, 99] as well as context-invariant perception [91].
In the neocortex, pyramidal neurons are thought to integrate feedforward and top-down, feedback information at their basal and apical dendrites, respectively [80, 101]. Notably, feedback has been proposed to modulate rather than drive neuronal activities [97]. This idea is supported by the unique morphological and physiological properties of pyramidal neurons, as activity in apical dendrites in vivo is generally not sufficient to drive somatic output responses alone [102, 79, 77, 73]. On the other hand, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity [78]. While the physiological properties of pyramidal neurons have been described extensively (see [101, ] for review), there is no consensus on the exact mechanism by which top-down signals from higher-order regions modulate somatic firing rates and update sensory representations.
Here, we propose a functional model that learns to use context only when such information is relevant, and ignores it otherwise, ultimately learning a useful mapping of top-down contextual information onto apical compartments. We examine character recognition tasks in cases where the characters are ambiguous and consider two different computational settings. First, the contextual representation is directly given to the model which learns to use this top-down input to resolve the ambiguities. Then, the representation of context is learned, from the temporal framing of the character in words or sequences of digits. In both settings, we show that the model learns to use apical modulation, driven by contextual inputs, to resolve sensory ambiguities. We find that a Hadamard (element-wise multiplicative) dendritic integration rule informed by physiological data enables strong performance in both settings of the task.
To gain insights into the learned mechanisms of top-down modulation involved in our model, we applied layer-wise relevance propagation [57] to identify the relative contributions of individual neurons and their apical compartments to network function. We identified a subset of neurons that are highly relevant for class-pairs which are selectively amplified by top-down signals arriving at their apical dendrites. We propose that, in principle, the representation of context can be learned independently from the impact that it has on ambiguity resolution. These findings provide a candidate mechanism for how neocortical pyramidal neurons integrate top-down contextual information at their apical dendrites to support robust perception in the face of ambiguous data.
2 Results
2.1 Functional model of context integration in apical dendrites
To model how apical dendrites integrate context to refine perceptual representations, we developed a task where correct classification of ambiguous characters requires appropriate integration of contextual cues (Fig. 1a). In this task, we trained a generative model to construct a dataset of ambiguous handwritten digits and letters based on MNIST and EMNIST, respectively. This generative model synthesizes images that are ambiguous between pairs of classes (see Methods). The resulting ambiguity was empirically quantified with a linear classifier trained on unambiguous images (fig. 1b). Inspired by anatomical features of pyramidal neurons, we next developed a model that implements distinct dendritic compartments, a basal compartment for integrating sensory information and an apical compartment for integrating contextual information (Fig. 1c). In the model, a population of $n$ pyramidal neurons has two sources of inputs, represented as vectors: a set of basal inputs, $\mathbf{b}=[b_{1},\ldots,b_{n}]$, representing the sensory stream, and a set of apical inputs, $\mathbf{a}=[a_{1},\ldots,a_{n}]$, representing the top-down stream. Input images, $\mathbf{x}$, are encoded by a pretrained network $f$, with parameters $\Theta_{b}$. $f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pretrained). The output of $f(x)$ is the basal vector $\mathbf{b}$.
The firing rate of pyramidal neurons, $\mathbf{h}=[h_{1},\ldots,h_{n}]$, is then determined by a thresholded element-wise (neuron-wise) combination of the basal and apical input vectors:
$\mathbf{h}=\sigma(\mathbf{b})\odot(\sigma(\mathbf{a})+1),$ (1)
where $\odot$ is an element-wise multiplication and $\sigma$ is the rectified linear unit (ReLU) activation function. We refer to this integration rule as ’Hadamard’. As can be seen from equation 1, the impact of the apical input vector, $\mathbf{a}$, on the firing rate, $\mathbf{h}$, is determined by $\sigma(\mathbf{a})+1$, which has elements that are by design always greater than or equal to one. As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity. This modulated basal activity is ultimately the firing rate of the modelled neurons, in-line with previous experimental findings [105]. Under these constraints, apical inputs affect neuronal activity only under the condition that they exceed their activation threshold (when $a>0$), similar to the requirement for regenerative calcium spikes for apical inputs to influence somatic spiking in real pyramidal neurons [79]. The output of the model is read-out from a feedforward linear transformation of the pyramidal neurons firing rates:
$\mathbf{\mu}=\mathbf{U}\mathbf{h},$ (2)
where
$\mathbf{U}$
is a pretrained linear transformation and $\mu$ represents a higher-order neuronal population. To obtain the apical inputs $\mathbf{a}$, the context, $\mathbf{c}$, (which is either derived from temporal context, or given directly, see below) is merged with the output and mapped onto the apical compartment using a multilayer perceptron (MLP), $g(\cdot)$, as follows:
$\mathbf{a}=g(\mathbf{c}\oplus\mathbf{U\sigma{(b)}}),$ (3)
where $\oplus$ is the concatenation operation, $\sigma$ is the ReLU activation, and $g$ is a 2-layer MLP (1 hidden layer) with learnable parameters denoted by $\Theta_{a}$. The model must use the combination of context and input to determine the representation of both ambiguous and unambiguous characters. For ambiguous characters, which are a mixture of two categories, the readout of the representations of the model will yield two plausible interpretations (one corresponding to the target category, and one corresponding to the contradictory category). The contradictory category refers to the class that induces the most confounding readout classification. Categories other than the target or contradictory we refer to as irrelevant categories. To test whether the model can learn useful apical modulations, we trained the synaptic weights $\Theta_{a}$ with the objective $L$ of modulating basal representations to match the target latent representations under various scenarios of input and context by gradient descent; Fig. 1e; pseudocode 1, and 5) This setup ensures that the top-down model is trained on both unambiguous and ambiguous samples, and that the model must learn to provide the appropriate modulatory signal only when it is relevant, ensuring that our model could learn to ignore contextual information when appropriate. This implementation reflects the fact that, in the mammalian brain, top-down modulation is available from multiple sources and present regardless of whether the sensory inputs are ambiguous or not, and whether the contextual sources are helpful for resolving ambiguity or not.
This suggests an approach where the problem of contextual integration is broken into two parts: (1) how apical dendrites locally learn to use contextual representations $c$ to solve the task and (2) how those contextual representations $c$ can be learned from the temporal frame.
For the first part, our initial experiments assume that context, $c$, is given in the form of a one-hot vector, a scenario we refer to as the ’oracle context’ (see Methods). Subsequent experiments for the second part address the problem of learning a representation of context, $c$, taking inspiration from the neocortex, where context is, in part, represented by higher-order regions with broader windows of temporal integration [65] (Fig. 1c; see Methods).
Figure 1: Functional model of context integration in apical dendrites. a, in our experimental paradigm, ambiguous characters (in yellow, top panels) are resolved by combining context signals (bottom panels). b, examples of unambiguous pairs and resulting ambiguous characters generated using EMNIST (left) and MNIST (right). c, model of pyramidal neurons with distinct dendritic compartments trained to solve a task of perceptual ambiguity. Before top-down modulation, two predictions have equal probability. After top-down modulation, context contributes to favoring one unambiguous representation. d, model implementation of dendritic integration using a hierarchical convolutional neural network. e, distinct combinations of input and context used to train the top-down network $g$. f, possible combinations of input and context used during training and analysis.
2.2 Resolving ambiguity in the oracle context
To first test if our model can categorize digits under contextual input represented as a one-hot vector (oracle context), we trained $g$ on the ambiguous MNIST/EMNIST task with a loss including all 5 scenarios of stimulus and context (see Methods). We assessed performance of the model in each scenario as the held-out test set accuracy on the pretrained category readout. and found model performance was significantly modulated by top-down signals (one-way ANOVA; $F_{4}$ = 10076.598; p $<$ 0.0001; n = 313 independent samples). Test set accuracy on ambiguous digits was $44.619\pm 8.073\%$ without top-down modulation, and became significantly higher with top-down modulation ($99.491\pm 1.306\%$; Tukey’s test; p $<$ 0.0001) if context was matching, but did not improve performance if context was irrelevant ($43.920\pm 8.160$; Tukey’s test; p = 0.4651; n = 313 independent replicates; Fig. 2a; Table 1).
[…]
Figure 2: Ambiguity can be resolved only when integrating top-down signals. a, readout test set accuracy before and after combining top-down signals over training epochs for all input-context combinations. Dashed line corresponds to chance level. b, low dimensional projection (t-SNE) of $h$ for unambiguous inputs (top), ambiguous inputs before (bottom left), and after (bottom right) top-down modulation. c, silhouette scores for $h$, for each class pair and for unambiguous inputs (left), ambiguous inputs before (center), and after (right) top-down modulation. d, cosine distance between $h$ in unambiguous conditions and basal, apical compartments and firing rate in each scenario.
[…]
Figure 3: Role of integration rules on ambiguity resolution. a, comparison between Hadamard and arithmetic sum integration rules. b, test set accuracy relative to contextual certainty using either integration rule in conditions where inputs are ambiguous and context is useful. c, same but in conditions where inputs are unambiguous and context contradicts the true label.
2.3 Single neuron mechanisms of top-down modulation
We next sought to understand the computational role played by top-down modulation of pyramidal apical dendrites in solving this contextual integration task. To this end, we developed a method to identify neurons relevant for solving the task. We then analysed their activations across the various contextualized input conditions to reveal the differential role played by their apical compartments. First, to define the sub-populations of pyramidal neurons performing specific computations, we employed layer-wise relevance propagation (LRP; [57]; supplementary Fig. 1a), a method typically used to explain which regions of an input contribute most to model output. Specifically, each neuron receives a score for its contribution to the model output for each image (Supplementary Fig. 1b). We then defined the subset of neurons most relevant for identifying a given class, $S_{C}$, as the minimum set of neurons required to account for 95% of total neuronal relevance when processing images from that class (Fig.  1c,d, see Methods). Sets of relevant neurons defined in this way exhibit low overlap between classes (Supplementary Fig. 1b-d, see Methods). We observed similar representational separability between classes for other scenario and given appropriate top-down inputs (Supplementary Fig. 1e). Additionally, these sets of relevant neurons are typically sparse and represent $\sim$15of the total neuronal population(Supplementary Fig. 1b,c).
[…]
To evaluate this hypothesis, we applied a mask on selected neurons at inference time and assessed the model’s performance. Specifically, we assess readout test set accuracy when inputs are ambiguous and context is helpful, and compare cases where class-pair overlapping neurons (as identified with LRP) are specifically masked (by setting their value to 0) to cases where other randomly selected neurons are masked. We also evaluate the effect of mask size and found that masking class-pair overlapping neurons specifically lead to significantly more rapid degradation of accuracy with increasing mask size compared to other neurons (Fig. 4d). This highlights the key role of these neurons’ apical compartments in integrating relevant contextual signals.
Figure 4: Class-pair overlapping neurons receive the strongest top-down signals. a, using layer-wise relevance propagation, we identify neurons most relevant for either of two classes (red and blue), or overlap of both classes (magenta). b, amplitude of the apical compartment (normalized over all neurons) for each input/context condition. Three distinct classes of neurons are compared: neurons most relevant for unambiguous class A (blue), B (red), or neurons most relevant for both A and B classes (magenta). c, we hypothesize that class-pair overlapping neurons receive the highest amplitude of top-down signals when processing ambiguous images, especially given matching context.
2.4 Deriving context from temporal information
Contextual priors can also be extracted from the temporal domain, specifically by leveraging past information to decode incoming ambiguous sensory inputs. To extend our model to cases where contextual signals are derived from the temporal domain, we trained a Gated Recurrent Unit (GRU) network to predict the arithmetic sum of two unambiguous digits as encoded by our pretrained feedforward weights $\mathbf{\Theta_{b}}$. The final output state of the GRU was then given as a contextual signal for our model. (fig. 5a; see Methods). We find that in cases where the input is ambiguous and the temporal sequence sums to a plausible interpretation of the ambiguous image, the model updates perceptual representations using contextual signals to successfully resolve ambiguities (fig. 5b, top). Importantly, context is effectively ignored in cases where input is unambiguous and contextual signals irrelevant (fig. 5b, bottom). Using low-dimensional projections of latent representations, we find that top-down context derived from the temporal domain effectively disentangled overlapping representations when inputs are ambiguous, similarly to the oracle case. (fig. 5c)
Task context ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra MNIST oracle 0.996 $\pm$ 0.000 0.983 $\pm$ 0.001 0.979 $\pm$ 0.002 0.461 $\pm$ 0.003 0.964 $\pm$ 0.001 MNIST temporal 0.963 $\pm$ 0.001 0.999 $\pm$ 0.000 0.991 $\pm$ 0.001 0.487 $\pm$ 0.005 0.990 $\pm$ 0.001 EMNIST oracle 0.993 $\pm$ 0.001 0.969 $\pm$ 0.003 0.936 $\pm$ 0.003 0.469 $\pm$ 0.004 0.903 $\pm$ 0.002
Table 1: Readout test set accuracy on MNIST and EMNIST with different context representations. Results are expressed for each scenario as mean % $\pm$ standard deviation.
Figure 5: Deriving context from temporal information. a, rationale for training a model that leverages context derived from temporal sequences through a recurrent network (GRU). b, example sequence used to generate context signals (as the arithmetic sum of two digits, left and center columns) to resolve ambiguity of an input image (right column). c, t-SNE projection of latent representation associated with ambiguous inputs before (left) and after (right) combining top-down signals.
Table 2: Ablation experiments on EMNIST for the integration rule. Context only $=0$ when the topdown network has access to feedforward latent $\mathbf{Ub}$ and $c$ (standard case), and context only $=1$ when only $c$ is provided as input. Three modulation (integration) rules are tested: hadamard (ours), sum, and concat (see 3. Concat refers to a concatenation of ’topdown’ weights on the basal weight matrix, akin to adding dendritic branches on the basal compartment.
ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra modulation context_only
0.989 $\pm$ 0.002 0.959 $\pm$ 0.003 0.931 $\pm$ 0.004 0.469 $\pm$ 0.003 0.841 $\pm$ 0.004 concat 0
0.794 $\pm$ 0.005 0.967 $\pm$ 0.003 0.938 $\pm$ 0.003 0.473 $\pm$ 0.004 0.903 $\pm$ 0.014 concat 1
0.993 $\pm$ 0.001 0.969 $\pm$ 0.003 0.936 $\pm$ 0.003 0.469 $\pm$ 0.004 0.903 $\pm$ 0.001 hadamard 0
0.913 $\pm$ 0.008 0.929 $\pm$ 0.004 0.901 $\pm$ 0.006 0.468 $\pm$ 0.005 0.738 $\pm$ 0.008 hadamard 1
0.994 $\pm$ 0.001 0.974 $\pm$ 0.001 0.944 $\pm$ 0.001 0.466 $\pm$ 0.007 0.920 $\pm$ 0.001 sum 0
0.792 $\pm$ 0.011 0.967 $\pm$ 0.005 0.937 $\pm$ 0.003 0.462 $\pm$ 0.013 0.904 $\pm$ 0.008 sum 1
3 Discussion
The exact mechanism by which top-down signals exert their functional role in pyramidal neurons remains unknown. One hypothesis is that, in the neocortex, apical dendrites of pyramidal neurons integrate top-down information from higher-order regions, which in turn modulate somatic firing rate for accurate and robust perception. A computational framework that respects known biophysical constraints while leveraging a task that requires integrating different types of context is ideal for testing this hypothesis. Here, we present (1) a functional model of contextual integration at pyramidal neurons that relies on the biophysical properties of apical and basal dendrites, and (2) a novel image classification task and dataset using a conditional generative model. Specifically for (1), we implement a Hadamard (multiplicative elementwise) integration rule based on known biophysical properties of apical dendrites which act as gain modulators of somatic activity [105, 80]. To solve (2), we train the top-down network (mapping of context onto apical dendrites) with the proposed loss function 5 by gradient descent, and assess the role of top-down signals in solving the task across 5 distinct input/context conditions. We find that the top-down network learns to refine ambiguous representations by applying sparse gain modulation to basal signals when contextual information is both available and relevant, but also learns to ignore contradictory or irrelevant contexts. This sparse selectivity is consistent with biased competition in the top-down modulation of pyramidal neurons. Our results show that it is in principle possible for local pyramidal circuits to learn a mapping of contextual representations onto their apical compartments to resolve perceptual ambiguity. Further analysis reveals that a sparse subset of class-pair overlapping neurons are preferentially gain-modulated by top-down contextual inputs, providing a candidate mechanism for how the model solves the perceptual ambiguities. This mixed selectivity could present distinct computational advantages across changing contexts [95].
3.1 Roles of top-down modulation
Top-down feedback is a prominent feature of the neocortical architecture for which several roles have been hypothesized. During learning, it has been proposed to act as a credit assignment signal [81, 70, 86, 59] and/or representing predictions [94, 93]. Besides its role in credit assignment, top-down feedback has also been proposed to support attentional processes [82] by dynamically highlighting the most relevant parts of an input, in both biological [100] and artificial systems [87], and could augment sensory processing with robustness to adversarial stimuli [61] or integration of contextual or task information [91, 106] . Given their biophysical properties, apical dendrites are particularly well poised to integrate these top-down contextual signals, and from a computational perspective, have been shown to be useful for multitask learning [106]. Our computational framework falls into the category of updating feedforward representations. We do not exclude other hypothesized roles for top-down feedback, however their unification under one framework is outside of the scope of this work.
Extending previous work on integration of topdown contextual signals via apical and basal dendrites, we focus specifically on a task in which the model must learn to differentially integrate relevant and irrelevant contextual signals at apical dendrites for accurate perception. This reflects the reality for brains, where not all contextual signals are useful or relevant to the task at hand.
3.2 Role of dendritic integration algorithms on function
Previous models have proposed to integrate feed-forward and feedback information streams using additive [87, 61] , multiplicative [91] rules, or both [106]. In [87, ], top-down signals come from past temporal context from higher-order modules, and are integrated with the feedforward stream via additive key-value attention ([103]). These works explore similar integration rules to ours, but the mathematical forms are not equivalent. Here we implement a Hadamard integration rule on the basis of a direct correspondence to biophysical principles. While the contextual integration task can be solved using different rules 2, under uncertainty, Hadamard integration enables the model to ignore contradictory contextual signals more efficiently compared to additive integration. Furthermore, as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike [106] and [91]. From an efficiency perspective, this eliminates the need to drive other neurons to fire, and uses less of the network capacity, at the cost of expressivity. It remains to be demonstrated whether our rule enables the model to perform well for more complex tasks.
3.3 Testable predictions
We find that our model can learn a non-linear mapping of top-down signals onto the apical compartment, a property that was recently described in pyramidal neurons of the visual cortex [67]. In addition to this compatibility with in vivo conditions, we propose several key links between experimental work and our model. Firstly, we found that under ambiguity, our model displayed a high diversity of neural responses as well as neurons that exhibit highly overlapping responses for specific class-pairs similar to in vivo reports that subsets of pyramidal neurons display a variety of responses given a mismatch between context and sensory evidence [74, 95]. Secondly, we find that masking apical compartments of class-pair overlapping neurons decreased task performance by a significant margin, up to nearly chance level. This is in contrast with in vivo experiments in non-human primates showing that inhibiting higher-order regions, and effectively shifting visual cortex representations to a strictly feed-forward regime, increased performance delays but did not prevent object recognition even for challenging images [75]. Finally, our model could learn a mapping of top-down feedback to the apical compartment independent of learning contextual representations which is compatible with high-order cortical regions, including the hippocampus, learning and updating contextual representations in absence of sensory inputs and yet play a major role in memory retrieval that necessitate contextual information [84].
One key testable prediction from our model is that neurons displaying mixed-selectivity would receive the strongest top-down signals at their apical dendrites. Testing this would require first conducting a controlled contextual integration task with an animal, and identifying a population from which a high degree of overlap is observed in neural responses for multiple categories or task variables. We hypothesize that silencing the apical compartments of these neurons would impair the ambiguity resolution ability of the animal.
3.4 Limitations and outlook
While our model combines functionality with key biophysical properties of dendrites, there are some important assumptions and simplifications that need to be considered. Firstly, our model is designed to implement some key features of cortical pyramidal neurons without focusing on a specific subpopulation that make up cortical layers, including layer 2/3 and layer 5 pyramidal neurons. Here, we assume that contextual signals are provided by a neurons in a higher-order region, which would likely correspond to layer 5 output, and integrated in apical dendrites of putative layer 2/3 pyramidal neurons [58, 96]. It is noteworthy though that layer 5 neurons can also send top-down projections to the basal compartment of layer 2/3 neurons, although to a lesser extent [72].
Secondly, we use a deep learning framework that does not respect Dale’s principle: namely, synaptic weights can be either negative and positive, giving the ability of presynaptic neurons to be both excitatory and inhibitory at the same time for distinct postsynaptic neurons. Although our model ignores this principle, we apply normalization across inputs, which could in principle mimic a role of the inhibitory neurons if they are non-selective. Additionally, Dale’s principle could theoretically be implemented to our deep learning framework with little sacrifice in performance [62].
Our model heavily abstracts temporal dynamics that are normally observed in biological circuits. Specifically, perception is associated with ever evolving neurophysiological responses that anchor to discrete states of sensory experience. For example, in the presence of ambiguous stimuli, human subjects can oscillate between two interpretations, a phenomenon termed ’perceptual bistability’ [71]. During perceptual bistability, neural activity alternates between discrete states [54, 55], following principles of continuous attractor dynamics [63]. In the future, our model could support more realistic temporal dynamics, which would enable the testing of a rich array of hypotheses regarding contextual integration.
In the future, we propose that our model could be extended to hierarchies of sensory modalities interacting with each other. Deep learning models that implement multi-modal interactions are an active area of research [92, 56, 89]. In these cases, representations from each modality are mapped to a joint high dimensional space and optimized via a contrastive self-supervised training objective. These models do not have the explicit objective to resolve conflicting or contradictory information from other modalities. In this context, our framework could be extended to multi-modal self-supervised learning for more robust perception in deep neural networks. Finally, such a model could be leveraged to test specific hypotheses of contextual integration in health and disease, and in particular in cases where the physiological properties of pyramidal neurons are known to be altered [76].
4 Methods
4.1 Ambiguous dataset generation
To create the dataset for our task, we first train a conditional variational autoencoder (CVAE) on standard MNIST/EMNIST [64, 98]. The purpose of the CVAE in this setup is to use the learned latent space to generate images that are ambiguous between two digits. While category-free interpolation in the latent space can be done with a vanilla VAE, the CVAE approach facilitates the generation of ambiguous images conditioned on the specified categories. We then concatenate the input to the decoder (latent $z$) with a linear combination of one-hot vectors representing the target digits. We define an image to be ambiguous when it lies at the $50\%\pm 5\%$ decision boundary between two classes of a classifier trained on standard MNIST/EMNIST. We collect a set of images in triplets, where two images are unambiguous from unique classes $y_{0},y_{1}$, and the third is ambiguous (using our criterion) between $y_{0},y_{1}$, generated by the CVAE. This dataset format simplifies the implementation of our task setup.
4.2 Model architecture
The model consists of three components: the pretrained backbone, the readout for classification, and the topdown network. We first train a variational autoencoder with a convolutional encoder and decoder to form smooth latent representations of unambiguous characters. Once the variational autoencoder backbone is trained, its weights are frozen and a readout (MLP) is trained to classify the latent representations.
4.2.1 Oracle context
In the oracle case, top-down signals are provided in the form of a one-hot vector:
$\mathbbm{1}_{y}(c_{i}):=\begin{cases}1,&\text{if $i=y$}.\\ 0,&\text{otherwise}\end{cases}$ (4)
$\forall i\in{1\dots C}$ where C is the number of classes, $|\mathbf{c}|=C$.
4.2.2 Temporal context
In the temporal case, context signals were provided by a Gated Recurrent Unit (GRU) network [60]. We trained a GRU with hidden state of size 128 to predict the modulo 10 sum of MNIST digits in a sequence of two digits’. The input to the GRU at time $t$ is not $x_{t}$ , but rather $\mathbf{Ub}_{t}$, encoded by the pretrained backbone. After processing a sequence of digits ($x_{t-2},x_{t-1}$), the final output state $o_{t}$ of the GRU was then used as a top-down context signal.
4.3 Training
4.3.1 Loss function
The topdown network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function composed of a sum over the following 5 scenarios: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra). All scenarios are therefore optimized together for every gradient update step. We use mean squared error as the loss criterion between the predicted latents ( computed as in equation 2) and the target latents for each scenario. Each minibatch is sampled from dataset $D$, and scenarios sampled from the minibatch and set of contexts $C$.
$\mathbf{L}(\Theta_{a};D)=\mathbbm{E}_{((x_{0},y_{0}),x_{ambig},(x_{1},y_{1}))\sim D,c_{match},c_{irrel}\sim C}[\sum_{s=1}^{5}||\mu_{s}^{*}-\hat{\mu_{s}}||_{2}^{2}]$ (5)
It should be noted that $x_{0},x_{1}$ are unambiguous images sampled from $D$ The optimum of $\Theta_{a}$ is given by:
$\Theta^{*}_{a}=argmin_{\Theta_{a}}L(\Theta_{a};D)$ (6)
Here, $\mu_{s}^{*}$ is the target representation, which is for each scenario and context:
$\mu_{s}^{*}:=\begin{cases}\mathbf{U}\mathbf{\sigma(b_{c_{match}})},&\text{if $s\in\{1\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})},&\text{if $s\in\{2,3,5\}$}\\ \mathbf{U}\mathbf{\sigma(b_{ambig})},&\text{if $s\in\{4\}$}\\ \end{cases}$ (7)
Where $b_{c}=f(x_{c}|\Theta_{b})$, and $b_{unambig}=f([x_{0},x_{1}]|\Theta_{b})$. Predicted latents are given by:
$\hat{\mu}_{s}:=\begin{cases}\mathbf{U}\mathbf{\sigma(b_{ambig})\odot(1+\sigma(a_{b_{ambig},c_{match}}))},&\text{if $s\in\{1\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})\odot(1+\sigma(a_{b_{unambig},c_{match}}})),&\text{if $s\in\{2\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})\odot(1+\sigma(a_{b_{unambig},c_{irrel}}})),&\text{if $s\in\{3\}$}\\ \mathbf{U}\mathbf{\sigma(b_{ambig})\odot(1+\sigma(a_{b_{ambig},c_{irrel}}))},&\text{if $s\in\{4\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})\odot(1+\sigma(a_{b_{unambig},c_{contra}}})),&\text{if $s\in\{5\}$}\\ \end{cases}$ (8)
Where $\mathbf{a_{b,c}}={g(\mathbf{Ub,c}|\Theta_{a})}$ (See pseudocode 1)
Algorithm 1 Training of the top-down model with oracle context
1:Dataset $D$, learning rate $\lambda$, dimension of basal activation $d_{b}$, number of classes $C$, top-down input $\mathbb{R}^{d_{b}+C}$, output: $\mathbb{R}^{d_{b}}$, bottom-up network (frozen) $f_{\Theta_{b}}$, latent projection (frozen) $\mathbf{U}$, top-down network $g(\dot{|}{\Theta_{a}})$, batch size B
2:
[…]
4:for triplet $\texttt{triplet minibatch}(\mathbf{x_{0}},\mathbf{y_{0}}),\mathbf{x_{ambig}},(\mathbf{x_{1}},\mathbf{y_{1}})\texttt{ sampled from }D$ do
5:     $y\leftarrow\{\mathbf{y_{0}},\mathbf{y_{1}}\}$
6:     $x\leftarrow\{\mathbf{x_{0},x_{1}}\}$
7:     # Concatenate all scenarios into single batch
8:     $\mathbf{c_{match}}\leftarrow random\_choice(y,size=B)$ Sample context uniformly from y over minibatch
[…]
10:     $\mathbf{c_{contra}}\leftarrow\mathbf{y_{1}}$
11:     # combine all vectors into one for efficiency
12:     $c\leftarrow cat(c_{match},c_{match},c_{irrel},c_{irrel},c_{contra})$
13:     $\mu^{*}\leftarrow cat(\mu_{c_{match}},\mu_{y},\mu_{y},\mu_{ambig},\mu_{y_{0}})$
14:     $x\leftarrow(x_{ambig},x,x,x_{ambig},x_{0})$
15:     $b\leftarrow f(x;\Theta_{b})$
16:     $a\leftarrow g(\mathbf{U}b\oplus c;\Theta_{a})$
17:     $\mathbf{\hat{h}}\leftarrow\textbf{relu}(b)\odot(1+\textbf{relu}(a))$
18:     $\mathbf{\hat{\mu}}\leftarrow\mathbf{U\hat{h}}$
19:     $Loss\leftarrow\textbf{mse}(\mathbf{\hat{\mu}},\mu^{*})$
20:     # autograd optimizer step
21:     $\Theta_{a}\leftarrow\Theta_{a}+\lambda\nabla_{\Theta_{a}}Loss$
4.3.2 Layer-wise Relevance Propagation Analysis
We applied layer-wise relevance propagation (LRP) to identify specific neurons whose activity, as described by $\mathbf{h}$ in our model, contributed most to the category readouts. For this analysis, we trained a readout to classify the standard categories (ie. digits 0-9) as well as all possible ambiguous pairs (ie. 3/5, 5/8, 1/7, etc.), which for MNIST gives a total number of categories $C=55$. To identify the subset of neurons that are most relevant for a given class $i$, we sorted neurons by highest to lowest average relevance per class. We then computed the normalized cumulative sum of relevance of neurons, and define $S_{i}$ as the minimum set of neurons required to obtained 95% of the total LRP for each class $i$ where $i\in{0,1,\cdots,C}$ (fig. 1d). We computed the separability between sets $S_{i}$, $S_{j}$ as $1-|S_{i}\cap S_{j}|/|S_{i}\cup S_{j}|$ (Fig. 1e).
4.4 Statistical analysis
Parametric tests where used when data distribution was normal and variance homogeneous, otherwise non-parametric tests were used and reported when appropriate. Meaning of abbreviated statistical terms. 1-ANOVA: one-way ANOVA; 2-ANOVA, two-way ANOVA; RM-ANOVA, repeated-measure ANOVA. Error bars and bands represent standard error of the mean, unless stated otherwise. *, p $<$ 0.05; **, p $<$ 0.01; ***, p $<$ 0.001; ****, p $<$ 0.0001; n.s., not significant.
4.5 Code availability
All experiments were carried out using PyTorch. We use the MNIST and EMNIST datasets, where number of output classes are $C=10$ for MNIST, and $C=26$ for EMNIST. The code for models and data analysis is publicly available under: https://github.com/ABL-Lab/expectation-clamp The code used to generate the ambiguous datasets as well as links to pre-generated datasets are available under: https://github.com/ABL-Lab/ambiguous-dataset
5 Acknowledgements
We thank Roberto Araya, Doina Precup, Irina Rish, and Yoshua Bengio for helpful discussions. This study was supported by funding from the Institute for Data Valorization (IVADO), the Unifying Neuroscience and Artificial Intelligence - Québec (UNIQUE) research center, the CHU Sainte-Justine Research Center (CHUSJRC), Fonds de Recherche du Québec–Santé (FRQS), the Canada CIFAR AI Chairs Program, the Quebec Institute for Artificial Intelligence (Mila), and Google. B.A.R. was supported by NSERC (Discovery Grant: RGPIN-2020-05105; Discovery Accelerator Supplement: RGPAS-2020-00031), CIFAR (Canada AI Chair; Learning in Machine and Brains Fellowship) and the Canada First Research Excellence Fund (CFREF Competition 2, 2015-2016) awarded to the Healthy Brains, Healthy Lives initiative at McGill University, through the Helmholtz International BigBrain Analytics and Learning Laboratory (HIBALL). Compute infrastructure was supported through a grant of computing time to E.B.M. from the Digital Research Alliance of Canada. N.I. received additional support from a NSERC Undergraduate student research award, and IVADO and UNIQUE excellence fellowships. M.T received additional support from the HBHL Graduate Fellowship. B.T.G. received additional support from a UNIQUE excellence fellowship.

### v2

@@
 Abstract
-One of the most striking features of neocortical anatomy is the presence of extensive top-down projections into primary sensory areas.
-Notably, many of these top-down projections impinge on the distal apical dendrites of pyramidal neurons, where they exert a modulatory effect, altering the gain of responses.
-It is thought that these top-down projections carry contextual information that can help animals to resolve ambiguities in sensory data.
-However, it has yet to be demonstrated how such modulatory connections to the distal apical dendrites can serve this computational function.
-Here, we develop a computational model of pyramidal cells that integrates contextual information from top-down projections to apical compartments with sensory representations driven by bottom-up projections to basal compartments.
-When input stimuli are ambiguous and relevant contextual information is available, the apical feedback modulates the basal signals to recover unambiguous sensory representations.
-Importantly, when stimuli are unambiguous, contextual information which is irrelevant or opposes sensory evidence is appropriately ignored by the model.
-By generalizing the task to temporal sequences, we further show that our model can learn to integrate contextual information across time.
-Using layer-wise relevance propagation, we extract the importance of individual neurons to the prediction of each category, revealing that neurons that are most relevant for the overlap of categories receive the largest magnitude of top-down signals, and are necessary for solving the task.
-This work thus provides a proof-of-concept demonstrating how the top-down modulatory inputs to apical dendrites in sensory regions could be used by the cortex to handle the ambiguities that animals encounter in the real world.
+One of the hallmark features of neocortical anatomy is the presence of extensive top-down projections into primary sensory areas, with many impinging on the distal apical dendrites of pyramidal neurons.
+While it is known that they exert a modulatory effect, altering the gain of responses, their functional role remains an active area of research.
+It is hypothesized that these top-down projections carry contextual information that can help animals to resolve ambiguities in sensory data.
+One proposed mechanism of contextual integration is a non-linear integration of distinct input streams at apical and basal dendrites of pyramidal neurons.
+Computationally, however, it is yet to be demonstrated how such an architecture could leverage distinct compartments for flexible contextual integration and sensory processing when both sensory and context signals can be unreliable.
+Here, we implement an augmented deep neural network with distinct apical and basal compartments that integrates a) contextual information from top-down projections to apical compartments, and b) sensory representations driven by bottom-up projections to basal compartments.
+In addition, we develop a new multi-scenario contextual integration task using a generative image modeling approach.
+In addition to generalizing previous contextual integration tasks, it better captures the diversity of scenarios where neither contextual nor sensory information are fully reliable.
+To solve this task, this model successfully learns to select among integration strategies.
+Specifically, when input stimuli and contextual information are contradictory, the performance of deep neural networks augmented with our "apical prior" exceeds that of single-compartment networks with otherwise equivalent architecture.
+We further show that our model can learn to integrate contextual information across time by using a task with input sequences and a gated recurrent unit layer to generate contextual signals.
+Using layerwise relevance propagation, we extract the relevance of individual neurons to the model’s predictions, revealing that a sparse subset of neurons encoding features of the context-relevant categories receive the largest magnitude of top-down signals when context and sensory input are aligned.
+We further show that this sparse gain modulation is necessary for best performance on the task.
+Altogether, this suggests that the "apical prior" and the biophysically-inspired non-linear integration rule could be key components necessary for handling the ambiguities that animals encounter in the diverse contexts of the real world.
 1 Introduction
-Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information [94, 88].
-Contextual information can be derived from a number of sources, including the spatial and temporal domains [65], task demands [69], as well as other sensory modalities.
+Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information (Rao and Ballard, 1999; Mumford, 1992).
+Contextual information can be derived from a number of sources, including the spatial and temporal domains (Eichenbaum, 2017), task demands (Gilbert and Li, 2013), as well as other sensory modalities.
 Several key neuroanatomical and cellular features could support the computations associated with context integration.
-In the mammalian brain, context has been proposed to be conveyed by top-down feedback pathways which are abundant in sensory regions of the neocortex [66, 85, 72].
-Perturbations of top-down pathways from higher order areas have been associated with delayed object recognition [75] as well as disrupted stimulus response curves in lower-order regions [104, 90], which may reflect deficits in contextual processing.
-Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images [68, 99] as well as context-invariant perception [91].
-In the neocortex, pyramidal neurons are thought to integrate feedforward and top-down, feedback information at their basal and apical dendrites, respectively [80, 101].
-Notably, feedback has been proposed to modulate rather than drive neuronal activities [97].
-This idea is supported by the unique morphological and physiological properties of pyramidal neurons, as activity in apical dendrites in vivo is generally not sufficient to drive somatic output responses alone [102, 79, 77, 73].
-On the other hand, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity [78].
-While the physiological properties of pyramidal neurons have been described extensively (see [101, ] for review), there is no consensus on the exact mechanism by which top-down signals from higher-order regions modulate somatic firing rates and update sensory representations.
-Here, we propose a functional model that learns to use context only when such information is relevant, and ignores it otherwise, ultimately learning a useful mapping of top-down contextual information onto apical compartments.
-We examine character recognition tasks in cases where the characters are ambiguous and consider two different computational settings.
-First, the contextual representation is directly given to the model which learns to use this top-down input to resolve the ambiguities.
-Then, the representation of context is learned, from the temporal framing of the character in words or sequences of digits.
-In both settings, we show that the model learns to use apical modulation, driven by contextual inputs, to resolve sensory ambiguities.
-We find that a Hadamard (element-wise multiplicative) dendritic integration rule informed by physiological data enables strong performance in both settings of the task.
-To gain insights into the learned mechanisms of top-down modulation involved in our model, we applied layer-wise relevance propagation [57] to identify the relative contributions of individual neurons and their apical compartments to network function.
-We identified a subset of neurons that are highly relevant for class-pairs which are selectively amplified by top-down signals arriving at their apical dendrites.
-We propose that, in principle, the representation of context can be learned independently from the impact that it has on ambiguity resolution.
+In the mammalian brain, context has been proposed to be conveyed by top-down feedback pathways which are abundant in sensory regions of the neocortex (Felleman and Van Essen, 1991; Markov et al., 2014; Harris et al., 2019).
+Perturbations of top-down pathways from higher order areas have been associated with delayed object recognition (Kar and DiCarlo, 2021) as well as disrupted stimulus response curves in lower-order regions (Wang et al., 2007; Nassi et al., 2013), which may reflect deficits in contextual processing.
+Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images (George et al., 2017; Spoerer et al., 2017) as well as context-invariant perception (Naumann et al., 2022), and context-dependent gating for multi-task learning Masse et al.
+(2018).
+In the neocortex, pyramidal neurons are thought to integrate feedforward and top-down, feedback information at their basal and apical dendrites, respectively (Larkum et al., 2007; Spruston, 2008).
+Notably, feedback has been proposed to modulate rather than drive neuronal activities (Sherman and Guillery, 1998).
+This idea is supported by the unique morphological and physiological properties of pyramidal neurons, as activity in apical dendrites in vivo is generally not sufficient to drive somatic output responses alone (Stuart and Spruston, 1998; Larkum et al., 1999; Larkum, 2013; Jarvis et al., 2018).
+However, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity (Larkum et al., 2004), serving as a potential stream that could carry contextual information.
+While the physiological properties of pyramidal neurons have been described extensively (see Spruston (2008) for review), there is no consensus on the exact mechanism by which top-down signals from higher-order regions modulate somatic firing rates and update sensory representations.
+Here, we propose a deep neural network architecture in which neurons are augmented specifically with apical and basal compartments and whose computation respects known biophysical properties of pyramidal neurons.
+We design a training paradigm such that, according to the context, the model updates its higher-level representations of sensory inputs, via the interaction of basal and apical compartments, in which the latter carries top-down contextual information.
+In tandem with the proposed framework, we develop a multi-scenario task with a dataset containing ambiguous images that are mixtures of two image categories, in which ambiguity can only be correctly resolved by integrating contextual information with sensory representations.
+We consider two versions of the problem setting.
+The first simplifies the contextual representation, which is given directly via a top-down oracle, and the model only needs to learn the apical parameters to correctly update the somatic activity.
+In the second, more challenging one, the model must use the temporal sequence as context and learn the appropriate contextual representations, along with the apical parameters to update somatic activity and ultimately solve the task.
+We show that the model learns to use apical modulation, driven by contextual inputs, to resolve sensory ambiguities.
+We find that a neuron-wise (element-wise) dendritic integration rule informed by physiological data enables strong performance in both settings of the task.
+To gain insights into the learned mechanisms of top-down modulation, we applied a standard method from the explainable AI literature ((Bach et al., 2015)) to identify the relative contributions of individual neurons and their apical compartments to network function.
+We identified a subset of neurons that are highly relevant for the matching category, and selectivelyamplified by top-down signals arriving at their apical dendrites.
+We propose that, in principle, 1) apical gain modulation of somatic activity and contextual integration are compatible, and is a simple yet efficient way to solve new tasks such as resolving perceptual ambiguity without altering what the model has previously learned, as it does not introduce or modify any synapses in the feed-forward pathway, and 2) that the apical gain modulatory signals and the representations of contextual information arriving at the aforementioned apical compartments can be decoupled.
+In a highly distributed system with local learning constraints such as the brain, the representation of context can be learned independently from the mapping of contextual signals to local populations to solve specific perceptual tasks.
 These findings provide a candidate mechanism for how neocortical pyramidal neurons integrate top-down contextual information at their apical dendrites to support robust perception in the face of ambiguous data.
@@
 2.1 Functional model of context integration in apical dendrites
-To model how apical dendrites integrate context to refine perceptual representations, we developed a task where correct classification of ambiguous characters requires appropriate integration of contextual cues (Fig.
+To model how apical dendrites integrate context to refine perceptual representations, we first developed a multi-scenario contextual integration task where correct classification of ambiguous characters requires appropriate integration of contextual cues (Fig.
 1a).
-In this task, we trained a generative model to construct a dataset of ambiguous handwritten digits and letters based on MNIST and EMNIST, respectively.
-This generative model synthesizes images that are ambiguous between pairs of classes (see Methods).
-The resulting ambiguity was empirically quantified with a linear classifier trained on unambiguous images (fig.
+To create the task, we trained a generative model on handwritten digits and letters (MNIST and EMNIST, respectively) to construct a dataset with highly ambiguous characters.
+We used the generative model to synthesize images that are a mixture of two classes (see Methods).
+The ambiguity of the synthesized images was empirically quantified as the proportion of the confusion of a linear classifier trained on unambiguous images (fig.
 1b).
@@
 In the model, a population of $n$ pyramidal neurons has two sources of inputs, represented as vectors: a set of basal inputs, $\mathbf{b}=[b_{1},\ldots,b_{n}]$, representing the sensory stream, and a set of apical inputs, $\mathbf{a}=[a_{1},\ldots,a_{n}]$, representing the top-down stream.
-Input images, $\mathbf{x}$, are encoded by a pretrained network $f$, with parameters $\Theta_{b}$.
-$f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pretrained).
+Input images, $\mathbf{x}$, are encoded by a pretrained network $f$, with parameters $\Theta_{b}$, where $f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pretrained).
 The output of $f(x)$ is the basal vector $\mathbf{b}$.
-The firing rate of pyramidal neurons, $\mathbf{h}=[h_{1},\ldots,h_{n}]$, is then determined by a thresholded element-wise (neuron-wise) combination of the basal and apical input vectors:
+The firing rate of pyramidal neurons, $\mathbf{h}=[h_{1},\ldots,h_{n}]$, is then determined by a thresholded neuron-wise (element-wise) combination of the basal and apical input vectors:
 $\mathbf{h}=\sigma(\mathbf{b})\odot(\sigma(\mathbf{a})+1),$ (1)
-where $\odot$ is an element-wise multiplication and $\sigma$ is the rectified linear unit (ReLU) activation function.
-We refer to this integration rule as ’Hadamard’.
-As can be seen from equation 1, the impact of the apical input vector, $\mathbf{a}$, on the firing rate, $\mathbf{h}$, is determined by $\sigma(\mathbf{a})+1$, which has elements that are by design always greater than or equal to one.
-As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity.
-This modulated basal activity is ultimately the firing rate of the modelled neurons, in-line with previous experimental findings [105].
-Under these constraints, apical inputs affect neuronal activity only under the condition that they exceed their activation threshold (when $a>0$), similar to the requirement for regenerative calcium spikes for apical inputs to influence somatic spiking in real pyramidal neurons [79].
-The output of the model is read-out from a feedforward linear transformation of the pyramidal neurons firing rates:
+where $\odot$ is an neuron-wise multiplication and $\sigma$ is the rectified linear unit (ReLU) activation function implementing basal and apical non-linearities, respectively.
+As can be seen from equation 1, the impact of the apical input vector, $\mathbf{a}$, on the firing rate, $\mathbf{h}$, is determined by $\sigma(\mathbf{a})+1$.
+As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity, in-line with previous experimental reports of a multiplicative role for apical inputs (Waters et al., 2003).
+Under these constraints, apical inputs affect neuronal activity only when they exceed their activation threshold (when $a>0$), analogous to observations that apical inputs influence somatic spiking only if they evoke a dendritic calcium spike (Larkum et al., 1999).
+To monitor the ability of the model to solve the perceptual task, the output of the model is read-out from a linear transformation of the pyramidal neurons firing rates:
 $\mathbf{\mu}=\mathbf{U}\mathbf{h},$ (2)
-where
-$\mathbf{U}$
-is a pretrained linear transformation and $\mu$ represents a higher-order neuronal population.
-To obtain the apical inputs $\mathbf{a}$, the context, $\mathbf{c}$, (which is either derived from temporal context, or given directly, see below) is merged with the output and mapped onto the apical compartment using a multilayer perceptron (MLP), $g(\cdot)$, as follows:
+where $\mu$ represents a higher-order neuronal population ($\mathbf{U}$ is a linear transformation to the latent space of the encoder $f$).
+To obtain the apical inputs $\mathbf{a}$, the context signal, $\mathbf{c}$, is concatenated with $\mu$, and mapped onto the apical compartment using a multilayer perceptron (MLP), $g(\cdot)$, as follows:
 $\mathbf{a}=g(\mathbf{c}\oplus\mathbf{U\sigma{(b)}}),$ (3)
-where $\oplus$ is the concatenation operation, $\sigma$ is the ReLU activation, and $g$ is a 2-layer MLP (1 hidden layer) with learnable parameters denoted by $\Theta_{a}$.
-The model must use the combination of context and input to determine the representation of both ambiguous and unambiguous characters.
-For ambiguous characters, which are a mixture of two categories, the readout of the representations of the model will yield two plausible interpretations (one corresponding to the target category, and one corresponding to the contradictory category).
-The contradictory category refers to the class that induces the most confounding readout classification.
-Categories other than the target or contradictory we refer to as irrelevant categories.
-To test whether the model can learn useful apical modulations, we trained the synaptic weights $\Theta_{a}$ with the objective $L$ of modulating basal representations to match the target latent representations under various scenarios of input and context by gradient descent; Fig.
-1e; pseudocode 1, and 5) This setup ensures that the top-down model is trained on both unambiguous and ambiguous samples, and that the model must learn to provide the appropriate modulatory signal only when it is relevant, ensuring that our model could learn to ignore contextual information when appropriate.
-This implementation reflects the fact that, in the mammalian brain, top-down modulation is available from multiple sources and present regardless of whether the sensory inputs are ambiguous or not, and whether the contextual sources are helpful for resolving ambiguity or not.
-This suggests an approach where the problem of contextual integration is broken into two parts: (1) how apical dendrites locally learn to use contextual representations $c$ to solve the task and (2) how those contextual representations $c$ can be learned from the temporal frame.
+where $\oplus$ is the concatenation operation, $\sigma$ is the ReLU activation, and $g$ is a MLP with 1 hidden layer, and its parameters are denoted by $\Theta_{a}$.
+As a population, these pyramidal neurons integrate sensory input character representations onto their basal dendritic compartment on one hand, and contextual information onto their apical dendritic compartment on the other hand.
+For ambiguous characters, which are a mixture of two categories, a readout of the representations of the model will yield two plausible interpretations, and contextual information is required to resolve the ambiguity.
+We consider the category which is compatible with the contextual input to be the target category.
+Conversely, a category which is a plausible interpretation of the input but not compatible with the contextual information is referred to as the contradictory category.
+We refer to all other categories as irrelevant categories.
+To test whether the model can learn apical modulations that are helpful for resolving ambiguous stimuli, we trained the synaptic weights $\Theta_{a}$ with the objective of modulating basal representations to match the target $\mu$ (representation after linear transform).
+We optimize this objective simultaneously for a set of 5 input-context scenarios (unambiguous or ambiguous input; helpful, contradictory or irrelevant context (Fig.
+1e) using a common gradient descent optimization algorithm (Adam, see pseudocode 1, and 5).
+The labels and corresponding descriptions we use for each scenario are provided explicitly below:
+  • ambig.Match: ambiguous input, helpful (matching) context
+  • unambig.Irrel: unambiguous input, irrelevant context
+  • ambig.Irrel: ambiguous input, irrelevant context
+  • unambig.Match: unambiguous input, helpful (matching) context
+  • unambig.Contra: unambiguous input, contradictory context
+This setup ensures that the top-down model is trained on both unambiguous and ambiguous examples, and that the model must learn to provide the appropriate modulatory signal only when contextual inputs provide useful information for solving the task, and ignore them when they are irrelevant or contradictory to sensory information.
+Our proposed implementation reflects the reality that, in the mammalian brain, top-down modulation is available from multiple sources regardless of the ambiguity of the sensory inputs or the relevance of the contextual information.
+This suggests an approach where the problem of contextual integration is broken into two parts: (1) how apical dendrites locally learn to use contextual representations $c$ to solve the multi-scenario task and (2) how useful contextual representations $c$ can be learned from available sources (e.g.
+temporal information).
 For the first part, our initial experiments assume that context, $c$, is given in the form of a one-hot vector, a scenario we refer to as the ’oracle context’ (see Methods).
-Subsequent experiments for the second part address the problem of learning a representation of context, $c$, taking inspiration from the neocortex, where context is, in part, represented by higher-order regions with broader windows of temporal integration [65] (Fig.
+Subsequent experiments for the second part address the problem of learning a representation of context, $c$, taking inspiration from the neocortex, where context is, in part, represented by higher-order regions with broader windows of temporal integration (Eichenbaum, 2017) (Fig.
 1c; see Methods).
@@
 b, examples of unambiguous pairs and resulting ambiguous characters generated using EMNIST (left) and MNIST (right).
-c, model of pyramidal neurons with distinct dendritic compartments trained to solve a task of perceptual ambiguity.
+c, model of pyramidal neurons with distinct dendritic compartments trained to solve a multi-scenario contextual integration task with ambiguity.
 Before top-down modulation, two predictions have equal probability.
 After top-down modulation, context contributes to favoring one unambiguous representation.
-d, model implementation of dendritic integration using a hierarchical convolutional neural network.
+d, implementation of dendritic integration in artificial neurons with apical and basal compartments.
+$f$ represents the feedforward mapping onto the basal compartment parameterized by the pretrained backbone, and $g$ the contextual top-down mapping onto the apical compartment, parameterized by its own neural network.
+The combination of the outputs $b$ and $a$ of $f$ and $g$, respectively, is applied neuron-wise via the multiplicative interaction $\sigma(b)*(1+\sigma(a))$.
 e, distinct combinations of input and context used to train the top-down network $g$.
@@
 2.2 Resolving ambiguity in the oracle context
-To first test if our model can categorize digits under contextual input represented as a one-hot vector (oracle context), we trained $g$ on the ambiguous MNIST/EMNIST task with a loss including all 5 scenarios of stimulus and context (see Methods).
-We assessed performance of the model in each scenario as the held-out test set accuracy on the pretrained category readout.
-and found model performance was significantly modulated by top-down signals (one-way ANOVA; $F_{4}$ = 10076.598; p $<$ 0.0001; n = 313 independent samples).
-Test set accuracy on ambiguous digits was $44.619\pm 8.073\%$ without top-down modulation, and became significantly higher with top-down modulation ($99.491\pm 1.306\%$; Tukey’s test; p $<$ 0.0001) if context was matching, but did not improve performance if context was irrelevant ($43.920\pm 8.160$; Tukey’s test; p = 0.4651; n = 313 independent replicates; Fig.
+To first test if our model can solve our multi-scenario task under contextual input represented as a one-hot vector (oracle context), we trained $g$ on the ambiguous MNIST/EMNIST task with a loss function that includes a loss term for each scenario We assessed performance of the model in each scenario as the accuracy of the readout on the test set, and found significant impact of top-down modulation.
+(one-way ANOVA; $F_{4}$ = 10076.598; p $<$ 0.0001; n = 313 independent samples).
+Specifically, test set accuracy on ambiguous digits was $46.3\pm 0.3\%$ without top-down modulation, and became significantly higher with top-down modulation ($98.5\pm 0.3\%$; Tukey’s test; p $<$ 0.0001) if context was matching, but did not significantly alter performance if context was irrelevant ($43.920\pm 8.160$; Tukey’s test; p = 0.4651; n = 313 independent replicates; Fig.
 2a; Table 2).
-[…]
-Figure 2: Ambiguity can be resolved only when integrating top-down signals.
+Importantly, the model did not learn to rely solely on top-down context, since performance remained high when the model was given contradictory context (Fig.
+2a; 2).
+To visualize the representations of the model on a population level, we projected unmodulated and modulated outputs $\mathbf{\mu}$ onto a two-dimensional manifold using t-SNE (Maaten (2008); Fig.
+2b).
+Ambiguous images whose representations exhibited high overlap between class-pairs were effectively disentangled after integrating matching contextual inputs (paired t-test, t99 = 6.0293, p $<$ 0.0001 for silhouette scores between ambiguous images before and after integrating top-down inputs; Fig.
+2c).
+These results show that our model learned to use contextual inputs provided to apical dendrites to modulate basal activity and resolve ambiguity.
+Figure 2: Our task is best solved when top-down signals are integrated correctly.
 a, readout test set accuracy before and after combining top-down signals over training epochs for all input-context combinations.
@@
 d, cosine distance between $h$ in unambiguous conditions and basal, apical compartments and firing rate in each scenario.
-[…]
+Task Context signal ambig (Baseline) ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra aMNIST oracle (one-hot) $46.3\pm 0.3$ $98.5\pm 0.3$ $98.5\pm 0.1$ $97.9\pm 0.1$ $46.0\pm 0.8$ $96.0\pm 0.4$ aEMNIST oracle (one-hot) $46.5\pm 0.5$ $94.2\pm 1.4$ $98.9\pm 0.2$ $96.2\pm 0.4$ $47.1\pm 0.8$ $94.9\pm 0.5$
+Table 1: Ambiguous MNIST and EMNIST (aMNIST, aEMNIST) test set classification accuracy across input-context scenarios with the hadamard integration rule.
+Results are expressed for each scenario as mean % $\pm$ standard deviation with 3 random seeds.
+Next, we evaluated the importance of having topdown apical compartments to handle contextual inputs compared with only having basal compartments (akin to point-wise spiking neurons, (Table.
+4).
+We found that, while there was not a large difference in the other 4 scenarios, there was an average increase in performance of 1.9% (MNIST) in the unambig.Contra scenario for the models which had apical compartments compared to those which only had basal compartments.
+Next, focusing our analysis on MNIST, we evaluated the importance of the integration rule which combines basal and apical activities4, 3a).
+We found a significant interaction between the integration rule and contextual certainty (2-ANOVA, F1 = 5.4741, p = 0.0346).
+When inputs are ambiguous and the context is highly certain (context vector with values $>$ 0.99 for the target class), we find that the additive and multiplicative (Hadamard) integration rules perform similarly.
+(pairwise t-test, t4 = 0.1708, p = 0.8678).
+For lower levels of certainty (context vector with values <=<=< = 0.6 for the target class), we note the following observations: 1) when context is matching, Hadamard integration is outperformed by Sum integration (pairwise t-test, t4 = -4.7677, p = 0.0089, and 2) when context is contradictory, Hadamard integration outperforms Sum integration (pairwise t-test, t4 = 8.9341, p = 0.0009; Fig.
+3b).
 Figure 3: Role of integration rules on ambiguity resolution.
-a, comparison between Hadamard and arithmetic sum integration rules.
-b, test set accuracy relative to contextual certainty using either integration rule in conditions where inputs are ambiguous and context is useful.
-c, same but in conditions where inputs are unambiguous and context contradicts the true label.
+a, comparison between Hadamard and additive integration rules.
+b, test set accuracy relative to contextual certainty using either integration rule in conditions where inputs are ambiguous and context is relevant (ambig.Match) or where inputs are unambiguous and context is contradictory to sensory input (unambig.Contra).
 2.3 Single neuron mechanisms of top-down modulation
-We next sought to understand the computational role played by top-down modulation of pyramidal apical dendrites in solving this contextual integration task.
-To this end, we developed a method to identify neurons relevant for solving the task.
-We then analysed their activations across the various contextualized input conditions to reveal the differential role played by their apical compartments.
-First, to define the sub-populations of pyramidal neurons performing specific computations, we employed layer-wise relevance propagation (LRP; [57]; supplementary Fig.
-1a), a method typically used to explain which regions of an input contribute most to model output.
-Specifically, each neuron receives a score for its contribution to the model output for each image (Supplementary Fig.
+We next sought to understand the computational role of top-down modulation of pyramidal apical dendrites in solving this contextual integration task.
+To this end, we used a method from the explainable AI literature to rank and identify the neurons most relevant for solving the task.
+We then compared their activations across the various context-input scenarios.
+First, to define the most relevant neurons, we employed layer-wise relevance propagation (LRP; (Bach et al., 2015); supplementary Fig.
+1a), a common method in explainable AI literature to explain, at each layer, the input’s contribution to the model output.
+Each neuron receives a score for its contribution to the model output for each image (Supplementary Fig.
 1b).
@@
 1b-d, see Methods).
-We observed similar representational separability between classes for other scenario and given appropriate top-down inputs (Supplementary Fig.
+We observed similar representational separability between classes for other scenarios and given appropriate top-down inputs (Supplementary Fig.
 1e).
@@
 1b,c).
-[…]
-To evaluate this hypothesis, we applied a mask on selected neurons at inference time and assessed the model’s performance.
-Specifically, we assess readout test set accuracy when inputs are ambiguous and context is helpful, and compare cases where class-pair overlapping neurons (as identified with LRP) are specifically masked (by setting their value to 0) to cases where other randomly selected neurons are masked.
-We also evaluate the effect of mask size and found that masking class-pair overlapping neurons specifically lead to significantly more rapid degradation of accuracy with increasing mask size compared to other neurons (Fig.
+It is noteworthy that our model is not explicitly provided with the ambiguous nature of input images, and therefore learns to extract this information to solve the task.
+In addition, the model learns to ignore top-down signals when the image is unambiguous and context is contradictory or irrelevant.
+Given that the apical compartment must combine contextual information and the contextually naive $\mu$ representation differentially based on input ambiguity, we hypothesized that solving the task effectively required the gain modulation of a subset of neurons encoding features specific to the relevant class.
+To assess this possibility, we registered the amplitude of apical signals arriving at LRP subsets pertaining to class-specific neurons and class-pair overlapping neurons in each scenario (Fig.
+4a-f).
+When comparing all input/context scenarios, we found that apical signals were highest when input images were ambiguous and context informative (1-ANOVA, F416=237.024, p $<$ 0.0001, for the main effect of input/context scenario; Fig.
+4b).
+In the ambig.Match scenario, neurons most relevant for the target class were associated with the highest amplitude in the apical compartment (1-ANOVA, F417=17.9445, p $<$ 0.0001, for the main effect of neuron class; Fig.
+4b).
+We hypothesized that target class specific neurons are also most modulated by top-down context, and that modulating these neurons is necessary to solve the perceptual ambiguity classification task.(Fig.
+4c).
+To evaluate this hypothesis, we applied a mask on selected neurons post-training, and assessed the model’s ability to solve the task.
+Specifically, we compare test set accuracy when inputs are ambiguous and context is relevant, under the condition where neurons (as identified with LRP) encoding features specific to the relevant class (which we refer to as context-relevant neurons) are masked out, compared to the condition where random neurons are masked.
+We also evaluate the effect of mask size and found that masking the context-relevant neurons specifically led to a steeper degradation of accuracy as we increase the number of masked neurons.
+(Fig.
 4d).
 This highlights the key role of these neurons’ apical compartments in integrating relevant contextual signals.
-Figure 4: Class-pair overlapping neurons receive the strongest top-down signals.
+Figure 4: Top-down has information about ambiguity, and in cases where inputs are ambiguous and context is aligned, resolves ambiguity by amplifying context-relevant neurons..
 a, using layer-wise relevance propagation, we identify neurons most relevant for either of two classes (red and blue), or overlap of both classes (magenta).
@@
 Three distinct classes of neurons are compared: neurons most relevant for unambiguous class A (blue), B (red), or neurons most relevant for both A and B classes (magenta).
-c, we hypothesize that class-pair overlapping neurons receive the highest amplitude of top-down signals when processing ambiguous images, especially given matching context.
+c, we hypothesize that context-relevant neurons receive the highest amplitude of top-down signals when processing ambiguous images, especially given matching context.
 2.4 Deriving context from temporal information
@@
 5c)
-Task context ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra MNIST oracle 0.996 $\pm$ 0.000 0.983 $\pm$ 0.001 0.979 $\pm$ 0.002 0.461 $\pm$ 0.003 0.964 $\pm$ 0.001 MNIST temporal 0.963 $\pm$ 0.001 0.999 $\pm$ 0.000 0.991 $\pm$ 0.001 0.487 $\pm$ 0.005 0.990 $\pm$ 0.001 EMNIST oracle 0.993 $\pm$ 0.001 0.969 $\pm$ 0.003 0.936 $\pm$ 0.003 0.469 $\pm$ 0.004 0.903 $\pm$ 0.002
-Table 1: Readout test set accuracy on MNIST and EMNIST with different context representations. Results are expressed for each scenario as mean % $\pm$ standard deviation.
+Task Context signal ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra seq-aMNIST temporal 96.3 $\pm$ 0.1 99.9 $\pm$ 0.0 99.1 $\pm$ 0.1 48.7 $\pm$ 0.5 99.0 $\pm$ 0.1
+Table 2: Sequential aMNIST (seq-aMNIST) test set classification accuracy across input-context scenarios with the hadamard integration rule.
+Results are expressed for each scenario as mean % $\pm$ standard deviation.
+In the temporal case, context signals are provided to the contextual integration network as GRU hidden state representations, conditioned on previous inputs in the sequence.
+Results are expressed for each scenario as mean % $\pm$ standard deviation with 3 random seeds.
 Figure 5: Deriving context from temporal information.
@@
 c, t-SNE projection of latent representation associated with ambiguous inputs before (left) and after (right) combining top-down signals.
-Table 2: Ablation experiments on EMNIST for the integration rule.
-Context only $=0$ when the topdown network has access to feedforward latent $\mathbf{Ub}$ and $c$ (standard case), and context only $=1$ when only $c$ is provided as input.
-Three modulation (integration) rules are tested: hadamard (ours), sum, and concat (see 3.
-Concat refers to a concatenation of ’topdown’ weights on the basal weight matrix, akin to adding dendritic branches on the basal compartment.
-ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra modulation context_only
-0.989 $\pm$ 0.002 0.959 $\pm$ 0.003 0.931 $\pm$ 0.004 0.469 $\pm$ 0.003 0.841 $\pm$ 0.004 concat 0
-0.794 $\pm$ 0.005 0.967 $\pm$ 0.003 0.938 $\pm$ 0.003 0.473 $\pm$ 0.004 0.903 $\pm$ 0.014 concat 1
-0.993 $\pm$ 0.001 0.969 $\pm$ 0.003 0.936 $\pm$ 0.003 0.469 $\pm$ 0.004 0.903 $\pm$ 0.001 hadamard 0
-0.913 $\pm$ 0.008 0.929 $\pm$ 0.004 0.901 $\pm$ 0.006 0.468 $\pm$ 0.005 0.738 $\pm$ 0.008 hadamard 1
-0.994 $\pm$ 0.001 0.974 $\pm$ 0.001 0.944 $\pm$ 0.001 0.466 $\pm$ 0.007 0.920 $\pm$ 0.001 sum 0
-0.792 $\pm$ 0.011 0.967 $\pm$ 0.005 0.937 $\pm$ 0.003 0.462 $\pm$ 0.013 0.904 $\pm$ 0.008 sum 1
 3 Discussion
-The exact mechanism by which top-down signals exert their functional role in pyramidal neurons remains unknown.
-One hypothesis is that, in the neocortex, apical dendrites of pyramidal neurons integrate top-down information from higher-order regions, which in turn modulate somatic firing rate for accurate and robust perception.
-A computational framework that respects known biophysical constraints while leveraging a task that requires integrating different types of context is ideal for testing this hypothesis.
-Here, we present (1) a functional model of contextual integration at pyramidal neurons that relies on the biophysical properties of apical and basal dendrites, and (2) a novel image classification task and dataset using a conditional generative model.
-Specifically for (1), we implement a Hadamard (multiplicative elementwise) integration rule based on known biophysical properties of apical dendrites which act as gain modulators of somatic activity [105, 80].
-To solve (2), we train the top-down network (mapping of context onto apical dendrites) with the proposed loss function 5 by gradient descent, and assess the role of top-down signals in solving the task across 5 distinct input/context conditions.
-We find that the top-down network learns to refine ambiguous representations by applying sparse gain modulation to basal signals when contextual information is both available and relevant, but also learns to ignore contradictory or irrelevant contexts.
-This sparse selectivity is consistent with biased competition in the top-down modulation of pyramidal neurons.
-Our results show that it is in principle possible for local pyramidal circuits to learn a mapping of contextual representations onto their apical compartments to resolve perceptual ambiguity.
-Further analysis reveals that a sparse subset of class-pair overlapping neurons are preferentially gain-modulated by top-down contextual inputs, providing a candidate mechanism for how the model solves the perceptual ambiguities.
-This mixed selectivity could present distinct computational advantages across changing contexts [95].
-3.1 Roles of top-down modulation
-Top-down feedback is a prominent feature of the neocortical architecture for which several roles have been hypothesized.
-During learning, it has been proposed to act as a credit assignment signal [81, 70, 86, 59] and/or representing predictions [94, 93].
-Besides its role in credit assignment, top-down feedback has also been proposed to support attentional processes [82] by dynamically highlighting the most relevant parts of an input, in both biological [100] and artificial systems [87], and could augment sensory processing with robustness to adversarial stimuli [61] or integration of contextual or task information [91, 106] .
-Given their biophysical properties, apical dendrites are particularly well poised to integrate these top-down contextual signals, and from a computational perspective, have been shown to be useful for multitask learning [106].
-Our computational framework falls into the category of updating feedforward representations.
-We do not exclude other hypothesized roles for top-down feedback, however their unification under one framework is outside of the scope of this work.
-Extending previous work on integration of topdown contextual signals via apical and basal dendrites, we focus specifically on a task in which the model must learn to differentially integrate relevant and irrelevant contextual signals at apical dendrites for accurate perception.
-This reflects the reality for brains, where not all contextual signals are useful or relevant to the task at hand.
-3.2 Role of dendritic integration algorithms on function
-Previous models have proposed to integrate feed-forward and feedback information streams using additive [87, 61] , multiplicative [91] rules, or both [106].
-In [87, ], top-down signals come from past temporal context from higher-order modules, and are integrated with the feedforward stream via additive key-value attention ([103]).
-These works explore similar integration rules to ours, but the mathematical forms are not equivalent.
-Here we implement a Hadamard integration rule on the basis of a direct correspondence to biophysical principles.
-While the contextual integration task can be solved using different rules 2, under uncertainty, Hadamard integration enables the model to ignore contradictory contextual signals more efficiently compared to additive integration.
-Furthermore, as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike [106] and [91].
-From an efficiency perspective, this eliminates the need to drive other neurons to fire, and uses less of the network capacity, at the cost of expressivity.
-It remains to be demonstrated whether our rule enables the model to perform well for more complex tasks.
-3.3 Testable predictions
-We find that our model can learn a non-linear mapping of top-down signals onto the apical compartment, a property that was recently described in pyramidal neurons of the visual cortex [67].
-In addition to this compatibility with in vivo conditions, we propose several key links between experimental work and our model.
-Firstly, we found that under ambiguity, our model displayed a high diversity of neural responses as well as neurons that exhibit highly overlapping responses for specific class-pairs similar to in vivo reports that subsets of pyramidal neurons display a variety of responses given a mismatch between context and sensory evidence [74, 95].
-Secondly, we find that masking apical compartments of class-pair overlapping neurons decreased task performance by a significant margin, up to nearly chance level.
-This is in contrast with in vivo experiments in non-human primates showing that inhibiting higher-order regions, and effectively shifting visual cortex representations to a strictly feed-forward regime, increased performance delays but did not prevent object recognition even for challenging images [75].
-Finally, our model could learn a mapping of top-down feedback to the apical compartment independent of learning contextual representations which is compatible with high-order cortical regions, including the hippocampus, learning and updating contextual representations in absence of sensory inputs and yet play a major role in memory retrieval that necessitate contextual information [84].
-One key testable prediction from our model is that neurons displaying mixed-selectivity would receive the strongest top-down signals at their apical dendrites.
-Testing this would require first conducting a controlled contextual integration task with an animal, and identifying a population from which a high degree of overlap is observed in neural responses for multiple categories or task variables.
-We hypothesize that silencing the apical compartments of these neurons would impair the ambiguity resolution ability of the animal.
-3.4 Limitations and outlook
-While our model combines functionality with key biophysical properties of dendrites, there are some important assumptions and simplifications that need to be considered.
+The ability to appropriately use and switch contexts to make sense of the vast stream of sensory information is an important feature of robust cognitive systems.
+It is thought that in the neocortex, contextual representations in higher order regions modulate lower sensory regions through top-down interactions arriving on the apical dendrites of pyramidal neurons.
+While it is known that apical dendrites modulate neuronal firing, their specific computational role in cognitive processes such as contextual integration remains poorly understood.
+Here, we tested whether the apical compartment represents an architectural bias that is functionally useful for contextual integration in neural networks.
+To this end, we implemented (1) a neural network architecture with distinct apical compartments and a neuron-wise integration rule based on the observation that apical dendrites are gain modulators of somatic activity (Waters et al., 2003; Larkum et al., 2007), and developed (2) an ambiguous image classification task that requires contextual integration to be solved.
+We trained (1) the contextual modulator (top-down synapses onto apical dendrites) by gradient descent to simultaneously learn all input-context conditions defined in our task.
+We found that this bio-inspired dendritic architecture outperformed a single-compartment model which processes both sensory and contextual information in the compartment.
+In this task setting, we find there is an inherent benefit to have contextual information propagated via a distinct apical compartment.
+By applying LRP, we find that the learned top-down modulation of pyramidal neurons is consistent with the principles of biased competition (Spratling, 2008).
+Specifically, by analyzing the apical amplitude on class-specific neurons for ambiguous stimuli, we found the highest gain modulation in class-specific neurons that matched the target class.
+Moreover, we found that the level of competition introduced by the context adapted to the level of ambiguity in the task: The top-down network learned to a) refine ambiguous representations by applying strong but sparse gain modulation to basal signals when contextual information is both available and relevant, and b) provide weak modulation in the case of contradictory or irrelevant contexts or unambiguous stimuli.
+3.1 Comparison with existing computational frameworks
+From a computational perspective, apical dendrites have been previously shown to support some types of contextual integration in biophysically realistic neurons and ANNs (Naumann et al., 2022; Wybo et al., 2022; George et al., 2020; Baronig and Legenstein, 2024).
+The tasks used in these previous works are multi-task, single scenario problem settings, meaning that context (e.g.
+one-hot encoded task id) is always assumed to be relevant.
+Alternative approaches to flexible contextual integration have involved explicitly learning to prevent associations between irrelevant context and sensory inputs ((Baronig and Legenstein, 2024)), which represents a successful example of how to handle spurious contexts in a simple task.
+As an extension of this approach, our model must infer how to handle the interaction between sensory input and context across distinct tasks and scenarios, where context is relevant, irrelevant, or contradictory.
+Crucially, we only found a significant advantage of the apical prior over other integration approaches in the scenario where inputs are unambiguous and context is contradictory (and should be ignored), with an accuracy increase of 1.9% (MNIST) for models with distinct apical compartments.
+This highlights a specific advantage for apical dendrites in the role of contextual integration, where contextual relevance is crucial to performance.
+Furthermore, while the task can be solved using either the Hadamard or additive rule (Appendix 4), we found that, under uncertainty, the Hadamard integration rule enabled the model to ignore contradictory contextual signals more effectively compared to additive integration.
+Another key difference with most previous work is that as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike Wybo et al.
+(2022) and (Naumann et al., 2022), but inline with experimental evidence (Larkum et al., 1999).
+From an efficiency perspective, this maintains the sparsity of the network.
+One notable prior work which applies context-dependent modulation to feedforward networks is context-dependent gating (XdG) (Masse et al., 2018), which could be comparable with our basal-only ablation as it is a feedforward architecture.
+The main difference with our basal-only ablation is that XdG a) is not informed by intermediate sensory representations, and as most prior works, assumes context always agrees with sensory input.
+Therefore, its is unlikely that XdG, implemented as-is, can deal with cases where context should be overridden by sensory input, as in the unambig.Contra case.
+The advantage of apical dendrites are clear in our multi-scenario contextual integration task.
+In order to compare our results with previous work, we examined the effect of our integration mechanism in a standard multi-task setting (given task-ids).
+Under these conditions we found that that all variants perform similarly well.
+This can be explained by the setting being simpler, whereas the distinct apical compartment shines under more varied and challenging types of contextual integration that require inference of ambiguity.
+While not explicitly modular, our framework could be seen as in accordance with the Global Workspace Theory along with the "information bottleneck" (to discard irrelevant information, which can include contextual information) Goyal et al.
+(2022).
+Under both frameworks, representations can be learned in a distributed and scalable manner, and mapped to local populations as context signals (in our case via apical compartments), and through biased competition, integrated with bottom-up representations to solve perceptual tasks.
+As highlighted in their work, this frees up the capacity of a learning system for more tasks, as it can more efficiently reuse what it has previously learned and identify task-relevant information, compared to single-compartment, single-integration models.
+3.2 Predictions: What is necessary for robust contextual integration?
+Until now, we have shown that in our contextual integration task, the apical prior provides a specific advantage compared to single-compartment integration when sensory or contextual signals are unreliable.
+Here, we aim to extend these findings to make broader predictions for what specific components/computations might be necessary for robust contextual integration, and how such a mechanism could plausibly be implemented in the cortex.
+Our results show that, given both sensory input and contextual input must be considered simultaneously, it may be sub-optimal to sum all activity together in the basal compartment, when sensory information and context are contradictory.
+At a higher level, this highlights a key difference between single and two-zone integration of information.
+As such, an interesting experiment would be to compare the single vs two-zone integration of bottomup representations and topdown contextual representations in simulations of neurons which more faithfully respect the biophysical constraints of pyramidal neurons in the neocortex, to observe whether our results hold in a more biologically plausible setting, or if it is strictly a feature of ANNs.
+Furthermore, based on our results, we can infer that the apical compartment encodes some information about the ambiguity of the sensory input, and the relevance of context for a given sensory input.
+An interesting experiment could involve silencing or deactivating apical dendrites in more realistic neurons (biophysical simulation or biological), and observing the effect on accurate perception in a similar contextual integration task.
+Mutual information between apical and basal activation is positive for the ambig.Match scenario, where we see a reduction in entropy of the prediction after observing the contextual signal.
+On the other hand, mutual information is near 0 for the unambig.Contra scenario, which aligns with our main findings (see Appendix, Table 3).
+This is not surprising given that we performed supervised training.
+We explicitly trained under the assumption that local representations, given mis-matching context, should be invariant to sensory ambiguity.
+This could be interpreted as being surprised by a stimulus, but not having informative or relevant context to reduce the ambiguity of the stimulus.
+This is important for robust contextual integration if one subscribes to the hypothesis that, most available contextual signals in rich, real-world sensory experience are distractors for any one particular task, and should be discarded.
+Following this, we predict that a model cannot learn the relevance of context with out encountering scenarios with irrelevant context during training.
+As an ablation study, we could remove the unambig.Irrel scenario from our loss function.
+We expect to observe a significant drop in accuracy on that scenario, reflecting the inability to assess contextual relevance.
+Extending this idea to unsupervised learning, it is interesting to think about how the model could identify contextual relevance without a supervised signal and be invariant to sensory ambiguity under unhelpful context.
+In an unsupervised setting this could arise from an information-maximization learning rule like InfoNCE, which uses a contrastive loss with positive and negative pairs defined via augmentation of the input (positive) against other inputs (negative).
+Unfortunately, the creation of positive and negative pairs relies heavily on domain-specific knowledge for useful data augmentations (eg.
+a rotation applied on an image).
+If it is possible to create contrastive pairs in a domain-agnostic manner, such an unsupervised learning framework could approximate the effects of supervised training shown in our results.
+Given the fact that most associations of context and input are either irrelevant or contradictory, and only a small subset of associations represent matching context, it is not surprising that the gain modulation is sparse.
+We expect also that our high-dimensional apical weight matrices could be well approximated with low-rank matrices without losing significant performance, in line with work in modulating deep learning models for downstream tasks via low-rank adapters Hu et al.
+(2021).
+While this method uses resources efficiently, it lacks the mechanism to separate irrelevant or contradictory from useful contexts that the apical prior excels at, which indicates a potential synergy between the two.
+We hypothesize the brain uses both mechanisms to learn to handle the large pool of available sources of context efficiently.
+A future direction could combine LORAs with the apical prior for contextual integration tasks.
+We find that our model could learn a mapping of top-down signals onto the apical compartment, which importantly, was learned separately from the contextual representations.
+This is compatible with high-order cortical regions such as the hippocampus, where the learning and updating of contextual representations is disentangled from how they modulate sensory representations (Maren et al., 2013).
+Because we assume the stability of the sensory and contextual representation, we cannot make any conclusions about how these two representations could be simultaneously co-learned.
+However, we speculate there should be some degree of stability in the sensory representations to learn to perform contextual integration.
+3.3 Limitations
+While our model unifies the function of contextual integration with key biophysical properties of dendrites, some key model assumptions and simplifications are worth discussing with respect to their empirical support.
 Firstly, our model is designed to implement some key features of cortical pyramidal neurons without focusing on a specific subpopulation that make up cortical layers, including layer 2/3 and layer 5 pyramidal neurons.
-Here, we assume that contextual signals are provided by a neurons in a higher-order region, which would likely correspond to layer 5 output, and integrated in apical dendrites of putative layer 2/3 pyramidal neurons [58, 96].
-It is noteworthy though that layer 5 neurons can also send top-down projections to the basal compartment of layer 2/3 neurons, although to a lesser extent [72].
-Secondly, we use a deep learning framework that does not respect Dale’s principle: namely, synaptic weights can be either negative and positive, giving the ability of presynaptic neurons to be both excitatory and inhibitory at the same time for distinct postsynaptic neurons.
-Although our model ignores this principle, we apply normalization across inputs, which could in principle mimic a role of the inhibitory neurons if they are non-selective.
-Additionally, Dale’s principle could theoretically be implemented to our deep learning framework with little sacrifice in performance [62].
+Here, we assume that contextual signals are provided by neurons in a higher-order region, which would likely correspond to layer 5 output, and integrated in apical dendrites of putative layer 2/3 pyramidal neurons (Bastos et al., 2012; Schuman et al., 2021).
+It is noteworthy though that layer 5 neurons can also send top-down projections to the basal compartment of layer 2/3 neurons, although to a lesser extent (Harris et al., 2019).
+Unlike other computational frameworks such as (George et al., 2020) which seek to incorporate the rich architectural diversity of the neocortex, we sought to simplify and abstract many features (using neural networks) in order to be able to train models to solve more challenging contextual integration tasks which could give useful insights that could not be seen with simpler tasks.
 Our model heavily abstracts temporal dynamics that are normally observed in biological circuits.
-Specifically, perception is associated with ever evolving neurophysiological responses that anchor to discrete states of sensory experience.
-For example, in the presence of ambiguous stimuli, human subjects can oscillate between two interpretations, a phenomenon termed ’perceptual bistability’ [71].
-During perceptual bistability, neural activity alternates between discrete states [54, 55], following principles of continuous attractor dynamics [63].
+For example, in the presence of ambiguous stimuli, human subjects can oscillate between two interpretations, a phenomenon termed ’perceptual bistability’ (Hardstone et al., 2021).
+During perceptual bistability, neural activity alternates between discrete states (Abeles et al., 1995; Akrami et al., 2009), following principles of continuous attractor dynamics (Daelli and Treves, 2010).
+3.4 Outlook
 In the future, our model could support more realistic temporal dynamics, which would enable the testing of a rich array of hypotheses regarding contextual integration.
 In the future, we propose that our model could be extended to hierarchies of sensory modalities interacting with each other.
-Deep learning models that implement multi-modal interactions are an active area of research [92, 56, 89].
-In these cases, representations from each modality are mapped to a joint high dimensional space and optimized via a contrastive self-supervised training objective.
-These models do not have the explicit objective to resolve conflicting or contradictory information from other modalities.
-In this context, our framework could be extended to multi-modal self-supervised learning for more robust perception in deep neural networks.
-Finally, such a model could be leveraged to test specific hypotheses of contextual integration in health and disease, and in particular in cases where the physiological properties of pyramidal neurons are known to be altered [76].
+Deep learning models that implement multi-modal interactions are an active area of research (Radford et al., 2021; Alayrac et al., 2022; Mustafa et al., 2022).
+While they are very good at multimodal generation with conditioning on one modality, they do not currently have the ability to resolve conflicting information and can be easily confused by contradictions in mixed-modality data.
+While these models employ transformer architectures that integrate context differently than MLPs, our loss function could inspire new approaches for robust multi-modal perception in deep neural networks.
+From a neurophysiology perspective, our framework could be leveraged to test specific hypotheses of contextual integration in health and disease, and in particular in cases where the physiological properties of pyramidal neurons are known to be altered (Kim et al., 2014).
 4 Methods
 4.1 Ambiguous dataset generation
-To create the dataset for our task, we first train a conditional variational autoencoder (CVAE) on standard MNIST/EMNIST [64, 98].
-The purpose of the CVAE in this setup is to use the learned latent space to generate images that are ambiguous between two digits.
-While category-free interpolation in the latent space can be done with a vanilla VAE, the CVAE approach facilitates the generation of ambiguous images conditioned on the specified categories.
-We then concatenate the input to the decoder (latent $z$) with a linear combination of one-hot vectors representing the target digits.
-We define an image to be ambiguous when it lies at the $50\%\pm 5\%$ decision boundary between two classes of a classifier trained on standard MNIST/EMNIST.
-We collect a set of images in triplets, where two images are unambiguous from unique classes $y_{0},y_{1}$, and the third is ambiguous (using our criterion) between $y_{0},y_{1}$, generated by the CVAE.
+We developed a new visual contextual integration task composed of multiple scenarios that incorporate varying levels of ambiguity of sensory input and informativeness of contextual information.
+To create the image dataset for our task, we used a digit-conditioned generative approach based on conditional variational autoencoders (CVAEs).
+First, we train a CVAE on standard a image dataset, either MNIST for digits or EMNIST for characters (Doersch, 2021; Sohn et al., 2015).
+We then use the learned latent space along with a digit-conditioned input vector to generate images that are ambiguous between two digits by interpolating between the one-hot vectors of the two digits.
+We then assess the ambiguous samples using a trained classifier on the original dataset (either MNIST or EMNIST), and discard images that are outside the $50\%(\pm 5\%)$ decision boundary between the two digits.
+We organize our dataset of images into triplets, where two images are unambiguous from digits $y_{0},y_{1}$, and the third is generated by the CVAE to be ambiguous between $y_{0},y_{1}$.
 This dataset format simplifies the implementation of our task setup.
-4.2 Model architecture
-The model consists of three components: the pretrained backbone, the readout for classification, and the topdown network.
-We first train a variational autoencoder with a convolutional encoder and decoder to form smooth latent representations of unambiguous characters.
-Once the variational autoencoder backbone is trained, its weights are frozen and a readout (MLP) is trained to classify the latent representations.
+4.2 Contextual integration model architecture
+The contextual integration model consists of three components: the pretrained backbone that computes representations of input images, the readout network for classifying their respective representations, and the topdown network which modulates the intermediate representations from the pretrained backbone based on contextual information.
+The backbone is implemented as a convolutional variational autoencoder (VAE), and was pretrained on unambiguous characters with the standard ELBO loss.
+VAEs are known to learn a smooth latent space, which we assume here to be an important ingredient to allow learning of top-down contextual modulations.
+Next, we freeze the VAE weights and train a readout (MLP) on the VAEs latent representations of images with a classification objective (cross entropy loss).
+Finally, we freeze these components and train the topdown network in two contextual tasks, with oracle and temporal context respectively, as described in the next sections.
+The topdown network $g(\cdot)$ is implemented as a multilayer perceptron (MLP) mapping the concatenated VAE latents and context vector to apical activations (eq.
+3), which then modulate the VAEs intermediate representation.
+Various modulation functions (hadamard, additive, concatenation) are explored in the main text.
 4.2.1 Oracle context
-In the oracle case, top-down signals are provided in the form of a one-hot vector:
+In the oracle case, contextsignals are provided to the top-down network $g$ in the form of a one-hot vector:
 $\mathbbm{1}_{y}(c_{i}):=\begin{cases}1,&\text{if $i=y$}.\\ 0,&\text{otherwise}\end{cases}$ (4)
@@
 4.2.2 Temporal context
-In the temporal case, context signals were provided by a Gated Recurrent Unit (GRU) network [60].
-We trained a GRU with hidden state of size 128 to predict the modulo 10 sum of MNIST digits in a sequence of two digits’.
-The input to the GRU at time $t$ is not $x_{t}$ , but rather $\mathbf{Ub}_{t}$, encoded by the pretrained backbone.
-After processing a sequence of digits ($x_{t-2},x_{t-1}$), the final output state $o_{t}$ of the GRU was then used as a top-down context signal.
+In the temporal case, the information is presented in a sequence, and preceding elements determine the contextual information available to the model at the current time $t$.
+Therefore, we leverage a Gated Recurrent Unit (GRU) network (Cho et al., 2014) to dynamically capture and represent the temporal dependencies in the sequence.
+We trained our GRU to predict the modulo 10 sum of MNIST digits in a sequence of two digits’, reflecting the task in our sequential MNIST/EMNIST dataset.
+An important detail is that the input to the GRU at time $t$ is not the image $x_{t}$ , but rather the latent $\mathbf{Ub}_{t}$, encoded by the pretrained backbone.
+After processing a sequence of digits ($x_{t-2},x_{t-1}$), the output state $o_{t}$ of the GRU was then used as a top-down context signal, in place of the one-hot context from the oracle case.
+We used a hidden state of size 128 for the GRU.
 4.3 Training
 4.3.1 Loss function
-The topdown network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function composed of a sum over the following 5 scenarios: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra).
-All scenarios are therefore optimized together for every gradient update step.
-We use mean squared error as the loss criterion between the predicted latents ( computed as in equation 2) and the target latents for each scenario.
-Each minibatch is sampled from dataset $D$, and scenarios sampled from the minibatch and set of contexts $C$.
+The topdown network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function jointly across the following 5 conditions: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra).
+We use mean squared error as the loss criterion between the predicted (computed as in equation 2) and the target latent representations.
+Each minibatch is sampled from dataset $D$, and training data for each condition are equally balanced in the minibatch.
 $\mathbf{L}(\Theta_{a};D)=\mathbbm{E}_{((x_{0},y_{0}),x_{ambig},(x_{1},y_{1}))\sim D,c_{match},c_{irrel}\sim C}[\sum_{s=1}^{5}||\mu_{s}^{*}-\hat{\mu_{s}}||_{2}^{2}]$ (5)
-It should be noted that $x_{0},x_{1}$ are unambiguous images sampled from $D$ The optimum of $\Theta_{a}$ is given by:
+$x_{0},x_{1}$ are unambiguous images sampled from $D$.
+The optimum of $\Theta_{a}$ is given by:
 $\Theta^{*}_{a}=argmin_{\Theta_{a}}L(\Theta_{a};D)$ (6)
 Here, $\mu_{s}^{*}$ is the target representation, which is for each scenario and context:
-$\mu_{s}^{*}:=\begin{cases}\mathbf{U}\mathbf{\sigma(b_{c_{match}})},&\text{if $s\in\{1\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})},&\text{if $s\in\{2,3,5\}$}\\ \mathbf{U}\mathbf{\sigma(b_{ambig})},&\text{if $s\in\{4\}$}\\ \end{cases}$ (7)
-Where $b_{c}=f(x_{c}|\Theta_{b})$, and $b_{unambig}=f([x_{0},x_{1}]|\Theta_{b})$. Predicted latents are given by:
+$\mu_{s}^{*}:=\begin{cases}\mathbf{U}\mathbf{\sigma(b_{c_{match}})},&\text{if $s\in\{1\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})},&\text{if $s\in\{2,3\}$}\\ \mathbf{U}\mathbf{\sigma(b_{ambig})},&\text{if $s\in\{4\}$}\\ \mathbf{U}\mathbf{\sigma(b_{c_{match}})},&\text{if $s\in\{5\}$}\end{cases}$ (7)
+Where $b_{c}=f(x_{c}|\Theta_{b})$ and
+$b_{unambig}=f([x_{0},x_{1}]|\Theta_{b})$.
+Predicted latents are given by:
 $\hat{\mu}_{s}:=\begin{cases}\mathbf{U}\mathbf{\sigma(b_{ambig})\odot(1+\sigma(a_{b_{ambig},c_{match}}))},&\text{if $s\in\{1\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})\odot(1+\sigma(a_{b_{unambig},c_{match}}})),&\text{if $s\in\{2\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})\odot(1+\sigma(a_{b_{unambig},c_{irrel}}})),&\text{if $s\in\{3\}$}\\ \mathbf{U}\mathbf{\sigma(b_{ambig})\odot(1+\sigma(a_{b_{ambig},c_{irrel}}))},&\text{if $s\in\{4\}$}\\ \mathbf{U}\mathbf{\sigma(b_{unambig})\odot(1+\sigma(a_{b_{unambig},c_{contra}}})),&\text{if $s\in\{5\}$}\\ \end{cases}$ (8)
-Where $\mathbf{a_{b,c}}={g(\mathbf{Ub,c}|\Theta_{a})}$ (See pseudocode 1)
+Where $\mathbf{a_{b,c}}={g(\mathbf{Ub,c}|\Theta_{a})}$.
 Algorithm 1 Training of the top-down model with oracle context
 1:Dataset $D$, learning rate $\lambda$, dimension of basal activation $d_{b}$, number of classes $C$, top-down input $\mathbb{R}^{d_{b}+C}$, output: $\mathbb{R}^{d_{b}}$, bottom-up network (frozen) $f_{\Theta_{b}}$, latent projection (frozen) $\mathbf{U}$, top-down network $g(\dot{|}{\Theta_{a}})$, batch size B
-2:
 […]
-4:for triplet $\texttt{triplet minibatch}(\mathbf{x_{0}},\mathbf{y_{0}}),\mathbf{x_{ambig}},(\mathbf{x_{1}},\mathbf{y_{1}})\texttt{ sampled from }D$ do
-5:     $y\leftarrow\{\mathbf{y_{0}},\mathbf{y_{1}}\}$
-6:     $x\leftarrow\{\mathbf{x_{0},x_{1}}\}$
-7:     # Concatenate all scenarios into single batch
-8:     $\mathbf{c_{match}}\leftarrow random\_choice(y,size=B)$ Sample context uniformly from y over minibatch
+3:for triplet $\texttt{triplet minibatch}(\mathbf{x_{0}},\mathbf{y_{0}}),\mathbf{x_{ambig}},(\mathbf{x_{1}},\mathbf{y_{1}})\texttt{ sampled from }D$ do
+4:     $y\leftarrow cat(\mathbf{y_{0}},\mathbf{y_{1}})$
+5:     $x\leftarrow cat(\mathbf{x_{0},x_{ambig},x_{1}})$
+6:     $\mathbf{c_{match}},\mathbf{c_{contra}}\leftarrow random\_choice(y,size=(2,B))\texttt{ \# Sample context from y without replacement}$
 […]
-10:     $\mathbf{c_{contra}}\leftarrow\mathbf{y_{1}}$
-11:     # combine all vectors into one for efficiency
-12:     $c\leftarrow cat(c_{match},c_{match},c_{irrel},c_{irrel},c_{contra})$
-13:     $\mu^{*}\leftarrow cat(\mu_{c_{match}},\mu_{y},\mu_{y},\mu_{ambig},\mu_{y_{0}})$
-14:     $x\leftarrow(x_{ambig},x,x,x_{ambig},x_{0})$
-15:     $b\leftarrow f(x;\Theta_{b})$
-16:     $a\leftarrow g(\mathbf{U}b\oplus c;\Theta_{a})$
-17:     $\mathbf{\hat{h}}\leftarrow\textbf{relu}(b)\odot(1+\textbf{relu}(a))$
-18:     $\mathbf{\hat{\mu}}\leftarrow\mathbf{U\hat{h}}$
-19:     $Loss\leftarrow\textbf{mse}(\mathbf{\hat{\mu}},\mu^{*})$
-20:     # autograd optimizer step
-21:     $\Theta_{a}\leftarrow\Theta_{a}+\lambda\nabla_{\Theta_{a}}Loss$
+8:     $c\leftarrow cat(c_{match},c_{irrel},c_{contra})$
+[…]
+10:     $\mu=\mathbf{U}b$
+11:     $\mu c\leftarrow\mu\oplus c$
+12:     $a\leftarrow g(\mu c;\Theta_{a})$
+[…]
+[…]
+15:     # autograd optimizer step
+16:     $\Theta_{a}\leftarrow\Theta_{a}+\lambda\nabla_{\Theta_{a}}Loss$
 4.3.2 Layer-wise Relevance Propagation Analysis

### v3

@@
 While it is known that they exert a modulatory effect, altering the gain of responses, their functional role remains an active area of research.
-It is hypothesized that these top-down projections carry contextual information that can help animals to resolve ambiguities in sensory data.
-One proposed mechanism of contextual integration is a non-linear integration of distinct input streams at apical and basal dendrites of pyramidal neurons.
+It is hypothesized that one of the roles of these top-down projections is to carry contextual information that can help animals to resolve ambiguities in sensory data.
+One proposed mechanism of contextual integration is a combination of input streams at distinct (separate integration zone) apical and basal dendrites of pyramidal neurons.
 Computationally, however, it is yet to be demonstrated how such an architecture could leverage distinct compartments for flexible contextual integration and sensory processing when both sensory and context signals can be unreliable.
@@
 1 Introduction
-Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information (Rao and Ballard, 1999; Mumford, 1992).
+Accurate perception relies on appropriate integration of context, as sensory signals contain incomplete or ambiguous information (Mumford, 1992; Rao and Ballard, 1999).
 Contextual information can be derived from a number of sources, including the spatial and temporal domains (Eichenbaum, 2017), task demands (Gilbert and Li, 2013), as well as other sensory modalities.
@@
 Perturbations of top-down pathways from higher order areas have been associated with delayed object recognition (Kar and DiCarlo, 2021) as well as disrupted stimulus response curves in lower-order regions (Wang et al., 2007; Nassi et al., 2013), which may reflect deficits in contextual processing.
-Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images (George et al., 2017; Spoerer et al., 2017) as well as context-invariant perception (Naumann et al., 2022), and context-dependent gating for multi-task learning Masse et al.
-(2018).
+Several feed-forward computational models of perception have implemented top-down pathways to enable object recognition of occluded images (George et al., 2017; Spoerer et al., 2017) as well as context-invariant perception (Naumann et al., 2022), and context-dependent gating for multi-task learning (Masse et al., 2018).
 In the neocortex, pyramidal neurons are thought to integrate feedforward and top-down, feedback information at their basal and apical dendrites, respectively (Larkum et al., 2007; Spruston, 2008).
@@
 This idea is supported by the unique morphological and physiological properties of pyramidal neurons, as activity in apical dendrites in vivo is generally not sufficient to drive somatic output responses alone (Stuart and Spruston, 1998; Larkum et al., 1999; Larkum, 2013; Jarvis et al., 2018).
-However, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity (Larkum et al., 2004), serving as a potential stream that could carry contextual information.
+However, activation of the apical compartment has been shown to act as a gain modulator, amplifying concurrent basal activity (Larkum et al., 2004), serving as a potential mechanism for contextual information.
 While the physiological properties of pyramidal neurons have been described extensively (see Spruston (2008) for review), there is no consensus on the exact mechanism by which top-down signals from higher-order regions modulate somatic firing rates and update sensory representations.
 Here, we propose a deep neural network architecture in which neurons are augmented specifically with apical and basal compartments and whose computation respects known biophysical properties of pyramidal neurons.
-We design a training paradigm such that, according to the context, the model updates its higher-level representations of sensory inputs, via the interaction of basal and apical compartments, in which the latter carries top-down contextual information.
+We design a training paradigm such that the model updates its representations of sensory inputs arriving at the basal compartment according to the top-down contextual information arriving at the apical compartment.
 In tandem with the proposed framework, we develop a multi-scenario task with a dataset containing ambiguous images that are mixtures of two image categories, in which ambiguity can only be correctly resolved by integrating contextual information with sensory representations.
@@
 We show that the model learns to use apical modulation, driven by contextual inputs, to resolve sensory ambiguities.
-We find that a neuron-wise (element-wise) dendritic integration rule informed by physiological data enables strong performance in both settings of the task.
 To gain insights into the learned mechanisms of top-down modulation, we applied a standard method from the explainable AI literature ((Bach et al., 2015)) to identify the relative contributions of individual neurons and their apical compartments to network function.
-We identified a subset of neurons that are highly relevant for the matching category, and selectivelyamplified by top-down signals arriving at their apical dendrites.
-We propose that, in principle, 1) apical gain modulation of somatic activity and contextual integration are compatible, and is a simple yet efficient way to solve new tasks such as resolving perceptual ambiguity without altering what the model has previously learned, as it does not introduce or modify any synapses in the feed-forward pathway, and 2) that the apical gain modulatory signals and the representations of contextual information arriving at the aforementioned apical compartments can be decoupled.
-In a highly distributed system with local learning constraints such as the brain, the representation of context can be learned independently from the mapping of contextual signals to local populations to solve specific perceptual tasks.
+We identified a subset of neurons that are highly relevant for the contextually-relevant category, and selectively amplified by top-down signals on apical dendrites.
+Our results show that apical gain modulation of somatic activity by contextual information is a simple yet efficient way to solve new tasks, such as resolving perceptual ambiguity.
+Moreover, the approach is flexible, because it is achieved without altering what the model has previously learned, as it does not introduce or modify any synapses in the feed-forward pathway.
 These findings provide a candidate mechanism for how neocortical pyramidal neurons integrate top-down contextual information at their apical dendrites to support robust perception in the face of ambiguous data.
@@
 We used the generative model to synthesize images that are a mixture of two classes (see Methods).
-The ambiguity of the synthesized images was empirically quantified as the proportion of the confusion of a linear classifier trained on unambiguous images (fig.
+The ambiguity of the synthesized images was empirically quantified via softmax on classifier predictions, where we select images near 50% for two classes (fig.
 1b).
@@
 In the model, a population of $n$ pyramidal neurons has two sources of inputs, represented as vectors: a set of basal inputs, $\mathbf{b}=[b_{1},\ldots,b_{n}]$, representing the sensory stream, and a set of apical inputs, $\mathbf{a}=[a_{1},\ldots,a_{n}]$, representing the top-down stream.
-Input images, $\mathbf{x}$, are encoded by a pretrained network $f$, with parameters $\Theta_{b}$, where $f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pretrained).
+Input images, $\mathbf{x}$, are encoded by a pre-trained network $f$, with parameters $\Theta_{b}$, where $f$ is a convolutional neural network reflecting initial stages of sensory processing (see Methods for more details on how the encoder is pre-trained).
 The output of $f(x)$ is the basal vector $\mathbf{b}$.
@@
 $\mathbf{h}=\sigma(\mathbf{b})\odot(\sigma(\mathbf{a})+1),$ (1)
-where $\odot$ is an neuron-wise multiplication and $\sigma$ is the rectified linear unit (ReLU) activation function implementing basal and apical non-linearities, respectively.
+where $\odot$ is an neuron-wise multiplication (Hadamard) and $\sigma$ is the rectified linear unit (ReLU) activation function implementing basal and apical non-linearities, respectively.
 As can be seen from equation 1, the impact of the apical input vector, $\mathbf{a}$, on the firing rate, $\mathbf{h}$, is determined by $\sigma(\mathbf{a})+1$.
-As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity, in-line with previous experimental reports of a multiplicative role for apical inputs (Waters et al., 2003).
-Under these constraints, apical inputs affect neuronal activity only when they exceed their activation threshold (when $a>0$), analogous to observations that apical inputs influence somatic spiking only if they evoke a dendritic calcium spike (Larkum et al., 1999).
-To monitor the ability of the model to solve the perceptual task, the output of the model is read-out from a linear transformation of the pyramidal neurons firing rates:
+As such, the apical activity serves as a thresholded neuron-wise gain modulator of the basal activity, in-line with previous experimental reports of a multiplicative role for apical inputs (Waters et al., 2003), an architectural bias of neocortical networks we refer to here as the "apical prior." Under these constraints, apical inputs affect neuronal activity only when they exceed their activation threshold (when $a>0$), analogous to observations that apical inputs influence somatic spiking only if they evoke a dendritic calcium spike (Larkum et al., 1999).
+Thus, this neuron-wise gain modulation of basal representations by the apical compartment via the Hadamard product is an implementation of the "apical prior."
+To monitor the ability of the model to solve the perceptual task, we read-out model predictions from a linear transformation of the pyramidal neurons firing rates:
 $\mathbf{\mu}=\mathbf{U}\mathbf{h},$ (2)
-where $\mu$ represents a higher-order neuronal population ($\mathbf{U}$ is a linear transformation to the latent space of the encoder $f$).
-To obtain the apical inputs $\mathbf{a}$, the context signal, $\mathbf{c}$, is concatenated with $\mu$, and mapped onto the apical compartment using a multilayer perceptron (MLP), $g(\cdot)$, as follows:
+where $\mu$ represents a higher-order neuronal population and $\mathbf{U}$ is a linear transformation (to the latent space of the encoder $f$).
+To obtain the apical inputs $\mathbf{a}$, the context signal, $\mathbf{c}$, is concatenated with $\mu$, and mapped onto the apical compartment using a multi-layer perceptron (MLP), $g(\cdot)$, as follows:
 $\mathbf{a}=g(\mathbf{c}\oplus\mathbf{U\sigma{(b)}}),$ (3)
 where $\oplus$ is the concatenation operation, $\sigma$ is the ReLU activation, and $g$ is a MLP with 1 hidden layer, and its parameters are denoted by $\Theta_{a}$.
-As a population, these pyramidal neurons integrate sensory input character representations onto their basal dendritic compartment on one hand, and contextual information onto their apical dendritic compartment on the other hand.
+As a population, these pyramidal neurons integrate sensory inputs (representations of characters) onto their basal dendritic compartment on one hand, and contextual information onto their apical dendritic compartment on the other hand.
 For ambiguous characters, which are a mixture of two categories, a readout of the representations of the model will yield two plausible interpretations, and contextual information is required to resolve the ambiguity.
-We consider the category which is compatible with the contextual input to be the target category.
+We consider the category which is compatible with the contextual input to be the target (match) category.
 Conversely, a category which is a plausible interpretation of the input but not compatible with the contextual information is referred to as the contradictory category.
 We refer to all other categories as irrelevant categories.
-To test whether the model can learn apical modulations that are helpful for resolving ambiguous stimuli, we trained the synaptic weights $\Theta_{a}$ with the objective of modulating basal representations to match the target $\mu$ (representation after linear transform).
-We optimize this objective simultaneously for a set of 5 input-context scenarios (unambiguous or ambiguous input; helpful, contradictory or irrelevant context (Fig.
-1e) using a common gradient descent optimization algorithm (Adam, see pseudocode 1, and 5).
+In total, we consider 5 possible scenarios.
 The labels and corresponding descriptions we use for each scenario are provided explicitly below:
@@
   • unambig.Contra: unambiguous input, contradictory context
+To test whether the contextual integration model incorporating the "apical prior" can learn apical modulations that are helpful for resolving ambiguous stimuli, we trained the synaptic weights $\Theta_{a}$ with the objective of modulating basal representations to match the target $\mu$ (representation after linear transform).
+We optimize this objective simultaneously for the set of all input-context scenarios listed above (Fig.
+1e and Eq.5) using a gradient descent optimization algorithm (Adam, see pseudocode 1).
 This setup ensures that the top-down model is trained on both unambiguous and ambiguous examples, and that the model must learn to provide the appropriate modulatory signal only when contextual inputs provide useful information for solving the task, and ignore them when they are irrelevant or contradictory to sensory information.
@@
 d, implementation of dendritic integration in artificial neurons with apical and basal compartments.
-$f$ represents the feedforward mapping onto the basal compartment parameterized by the pretrained backbone, and $g$ the contextual top-down mapping onto the apical compartment, parameterized by its own neural network.
+$f$ represents the feedforward mapping onto the basal compartment parameterized by the pre-trained backbone, and $g$ the contextual top-down mapping onto the apical compartment, parameterized by its own neural network.
 The combination of the outputs $b$ and $a$ of $f$ and $g$, respectively, is applied neuron-wise via the multiplicative interaction $\sigma(b)*(1+\sigma(a))$.
@@
 2a; 2).
-To visualize the representations of the model on a population level, we projected unmodulated and modulated outputs $\mathbf{\mu}$ onto a two-dimensional manifold using t-SNE (Maaten (2008); Fig.
+To visualize the representations of the model on a population level, we projected un-modulated and modulated outputs $\mathbf{\mu}$ onto a two-dimensional manifold using t-SNE (Maaten (2008); Fig.
 2b).
@@
 Figure 2: Our task is best solved when top-down signals are integrated correctly.
-a, readout test set accuracy before and after combining top-down signals over training epochs for all input-context combinations.
+a, readout test set accuracy without (black) and with (cyan) top-down signals for all input-context combinations.
 Dashed line corresponds to chance level.
-b, low dimensional projection (t-SNE) of $h$ for unambiguous inputs (top), ambiguous inputs before (bottom left), and after (bottom right) top-down modulation.
-c, silhouette scores for $h$, for each class pair and for unambiguous inputs (left), ambiguous inputs before (center), and after (right) top-down modulation.
-d, cosine distance between $h$ in unambiguous conditions and basal, apical compartments and firing rate in each scenario.
+b, low dimensional projection (t-SNE) of the latent representation, $\mu$, for unambiguous inputs (top), ambiguous inputs before (bottom left), and after (bottom right) top-down modulation.
+c, silhouette scores for latent representation, $\mu$, for each class pair and for unambiguous inputs (left), ambiguous inputs before (center), and after (right) top-down modulation.
+d, cosine distance between unambiguous activation, $h$, and basal, apical compartments and firing rate in each scenario.
 Task Context signal ambig (Baseline) ambig.Match unambig.Match unambig.Irrel ambig.Irrel unambig.Contra aMNIST oracle (one-hot) $46.3\pm 0.3$ $98.5\pm 0.3$ $98.5\pm 0.1$ $97.9\pm 0.1$ $46.0\pm 0.8$ $96.0\pm 0.4$ aEMNIST oracle (one-hot) $46.5\pm 0.5$ $94.2\pm 1.4$ $98.9\pm 0.2$ $96.2\pm 0.4$ $47.1\pm 0.8$ $94.9\pm 0.5$
@@
 Results are expressed for each scenario as mean % $\pm$ standard deviation with 3 random seeds.
-Next, we evaluated the importance of having topdown apical compartments to handle contextual inputs compared with only having basal compartments (akin to point-wise spiking neurons, (Table.
+Next, we evaluated the importance of having top-down apical compartments to handle contextual inputs compared with only having basal compartments (akin to point-wise spiking neurons, (Table.
 4).
 We found that, while there was not a large difference in the other 4 scenarios, there was an average increase in performance of 1.9% (MNIST) in the unambig.Contra scenario for the models which had apical compartments compared to those which only had basal compartments.
-Next, focusing our analysis on MNIST, we evaluated the importance of the integration rule which combines basal and apical activities4, 3a).
+Next, focusing our analysis on MNIST, we evaluated the importance of the integration rule which combines basal and apical activities (Fig. 3a, Appendix 1 Table. 4).
 We found a significant interaction between the integration rule and contextual certainty (2-ANOVA, F1 = 5.4741, p = 0.0346).
@@
 We then compared their activations across the various context-input scenarios.
-First, to define the most relevant neurons, we employed layer-wise relevance propagation (LRP; (Bach et al., 2015); supplementary Fig.
-1a), a common method in explainable AI literature to explain, at each layer, the input’s contribution to the model output.
-Each neuron receives a score for its contribution to the model output for each image (Supplementary Fig.
+First, to define the most relevant neurons, we employed layer-wise relevance propagation (LRP; (Bach et al., 2015); Appendix Fig.
+1a), a method in the explainable AI literature to reveal, at each layer, the input’s contribution to the model output.
+Each neuron receives a score for its contribution to the model output for each image (Appendix Fig.
 1b).
-We then defined the subset of neurons most relevant for identifying a given class, $S_{C}$, as the minimum set of neurons required to account for 95% of total neuronal relevance when processing images from that class (Fig.
+We then defined the subset of neurons most relevant for predicting a given class, e.g.
+class A or class B, as the minimum set of neurons required to account for 95% of total neuronal relevance when processing images from that class (Appendix Fig.
 1c,d, see Methods).
-Sets of relevant neurons defined in this way exhibit low overlap between classes (Supplementary Fig.
+Similarly, we define subsets of neurons most relevant for predicting that an image is ambiguous between two classes A and B (ambiguous AB).
+Sets of relevant neurons defined in this way exhibit low overlap between classes (Appendix Fig.
 1b-d, see Methods).
-We observed similar representational separability between classes for other scenarios and given appropriate top-down inputs (Supplementary Fig.
+We observed similar representational separability between classes for other scenarios (Appendix Fig.
 1e).
-Additionally, these sets of relevant neurons are typically sparse and represent $\sim$15of the total neuronal population (Supplementary Fig.
+Additionally, these sets of relevant neurons are typically sparse and represent $\sim$15of the total neuronal population (Appendix Fig.
 1b,c).
-It is noteworthy that our model is not explicitly provided with the ambiguous nature of input images, and therefore learns to extract this information to solve the task.
+It is noteworthy that our model is not explicitly provided with the ambiguous nature of the input images, and therefore learns to extract this information to solve the task.
 In addition, the model learns to ignore top-down signals when the image is unambiguous and context is contradictory or irrelevant.
-Given that the apical compartment must combine contextual information and the contextually naive $\mu$ representation differentially based on input ambiguity, we hypothesized that solving the task effectively required the gain modulation of a subset of neurons encoding features specific to the relevant class.
-To assess this possibility, we registered the amplitude of apical signals arriving at LRP subsets pertaining to class-specific neurons and class-pair overlapping neurons in each scenario (Fig.
+Given that the apical compartment must combine contextual information and the contextually naive $\mu$ representation differentially based on input ambiguity, we hypothesized that solving the task effectively required a high gain modulation specific to the subset of context-relevant neurons.
+To assess this possibility, we computed the average amplitude of apical signals arriving at LRP subsets pertaining to predicting each of two input classes (A,B) and ambiguous AB for each scenario of matching, irrelevant or contradictory context (Fig.
 4a-f).
@@
 4b).
-We hypothesized that target class specific neurons are also most modulated by top-down context, and that modulating these neurons is necessary to solve the perceptual ambiguity classification task.(Fig.
+Apical input onto neurons outside of these defined subpopulations ("other") remained low ($\sim$ 0.07-0.08) across all scenarios (data not shown), indicating there was no relationship between other neurons’ apical signal and the scenario.
+We also computed the basal amplitude for the same subsets of neurons defined above and found that there was a higher average amplitude on the matching relative to the contradictory subpopulation for unambiguous stimuli.
+(Fig.
 4c).
-To evaluate this hypothesis, we applied a mask on selected neurons post-training, and assessed the model’s ability to solve the task.
-Specifically, we compare test set accuracy when inputs are ambiguous and context is relevant, under the condition where neurons (as identified with LRP) encoding features specific to the relevant class (which we refer to as context-relevant neurons) are masked out, compared to the condition where random neurons are masked.
-We also evaluate the effect of mask size and found that masking the context-relevant neurons specifically led to a steeper degradation of accuracy as we increase the number of masked neurons.
-(Fig.
-4d).
+Likewise, there was a higher basal amplitude on the ambiguous subpopulation relative to other subpopulations for ambiguous inputs.
+Basal input onto the "other" subpopulation was found to be significantly lower for ambig.match and unambig.contra scenarios, but similar in magnitude to the unambig/ambig.irrel scenarios ($<$0.02, Fig.
+4c).
+These observations are indicative of a mechanism whereby ambiguous bottom-up representations are biased towards top-down representations by specific apical inputs Fig.
+4d, whereas unambiguous bottom-up representations are preserved due to non-specific top-down modulation Fig.
+4e.
+To evaluate if specific top-down modulation is necessary for solving the ambig.match task, we applied a mask on the apical inputs to the corresponding subpopulations post-training, and assessed the model’s ability to solve the task (Appendix 2-Fig.
+1).
+Specifically, we compare test set accuracy when inputs are ambiguous and context is relevant, under the condition where context-relevant apical inputs are masked out (activations set to 0), compared to the condition where random apical inputs are masked.
+We also evaluate the effect of mask size and found that masking the apical inputs to context-relevant neurons specifically led to a steeper degradation of accuracy as we increase the number of masked inputs.
 This highlights the key role of these neurons’ apical compartments in integrating relevant contextual signals.
-Figure 4: Top-down has information about ambiguity, and in cases where inputs are ambiguous and context is aligned, resolves ambiguity by amplifying context-relevant neurons..
-a, using layer-wise relevance propagation, we identify neurons most relevant for either of two classes (red and blue), or overlap of both classes (magenta).
+Figure 4: Top-down has information about ambiguity, and in cases where inputs are ambiguous and context is aligned, resolves ambiguity by amplifying context-relevant neurons.
+a, using layer-wise relevance propagation, we identify neurons most relevant for either of two classes (red and blue), or for ambiguous inputs between the two classes (magenta).
 b, amplitude of the apical compartment (normalized over all neurons) for each input/context condition.
-Three distinct classes of neurons are compared: neurons most relevant for unambiguous class A (blue), B (red), or neurons most relevant for both A and B classes (magenta).
-c, we hypothesize that context-relevant neurons receive the highest amplitude of top-down signals when processing ambiguous images, especially given matching context.
+Three distinct classes of neurons are compared: neurons most relevant for unambiguous class A (blue), B (red), or neurons most relevant for the ambiguous AB images (magenta).
+c, Normalized basal signal.
+left: in case of ambiguous stimuli, both plausible class-subpopulations are active along with the ambiguity detector subpopulation.
+right: in case of unambiguous stimuli, we observe a similar distribution, with the context-relevant neurons (class A) having higher average firing rate.
+d, a cartoon summary of the observed mechanism for the ambig.match-B case (ambiguous input, contextual input favouring class B), where class B neurons code for ’3’ (dark red) and class A neurons code for ’8’ (blue).
+Context-relevant neurons (dark red) receive the highest amplitude of top-down signals when processing ambiguous input.
+At basal dendrites, these sub-populations have low but non-zero firing rates, which allows them to be selectively amplified with top-down signals via the Hadamard integration rule, and thus dictate the output representation.
+e, for unambiguous input and contradictory context (unambig.contra), top-down input becomes non-specific and neurons are weakly amplified by the Hadamard rule.
+The output is then dictated by the neurons receiving the strongest basal input, in this case class B neurons (dark red).
+Output representations for unambiguous inputs are thus preserved when contextual inputs are added.
 2.4 Deriving context from temporal information
 Contextual priors can also be extracted from the temporal domain, specifically by leveraging past information to decode incoming ambiguous sensory inputs.
-To extend our model to cases where contextual signals are derived from the temporal domain, we trained a Gated Recurrent Unit (GRU) network to predict the arithmetic sum of two unambiguous digits as encoded by our pretrained feedforward weights $\mathbf{\Theta_{b}}$.
-The final output state of the GRU was then given as a contextual signal for our model.
+To extend our model to cases where contextual signals are derived from the temporal domain, we trained a Gated Recurrent Unit (GRU) network to predict the arithmetic sum of two unambiguous digits as encoded by our pre-trained feedforward weights $\mathbf{\Theta_{b}}$.
+The final output state of the GRU was then given as a contextual signal for our contextual integration model.
 (fig.
@@
 5b, top).
-Importantly, context is effectively ignored in cases where input is unambiguous and contextual signals irrelevant (fig.
+Importantly, context is effectively ignored in cases where the input is unambiguous and the contextual signals are irrelevant (fig.
 5b, bottom).
-Using low-dimensional projections of latent representations, we find that top-down context derived from the temporal domain effectively disentangled overlapping representations when inputs are ambiguous, similarly to the oracle case.
+Using low-dimensional projections of latent representations, we find that top-down context derived from the temporal domain effectively disentangled overlapping representations when inputs are ambiguous, similar to the oracle case.
 (fig.
@@
 3 Discussion
-The ability to appropriately use and switch contexts to make sense of the vast stream of sensory information is an important feature of robust cognitive systems.
-It is thought that in the neocortex, contextual representations in higher order regions modulate lower sensory regions through top-down interactions arriving on the apical dendrites of pyramidal neurons.
-While it is known that apical dendrites modulate neuronal firing, their specific computational role in cognitive processes such as contextual integration remains poorly understood.
-Here, we tested whether the apical compartment represents an architectural bias that is functionally useful for contextual integration in neural networks.
+The ability to appropriately use and switch contexts to make sense of a vast stream of sensory information is an important feature of robust cognitive systems.
+It is thought that in the neocortex, contextual representations in higher order regions modulate lower sensory regions through top-down interactions arriving on the apical dendrites of pyramidal neurons (Larkum, 2013; Phillips, 2017), an architectual bias we refer to here as the "apical prior." While it is known that apical dendrites modulate neuronal firing, their specific computational role in cognitive processes such as contextual integration remains poorly understood.
+Here, we tested whether the "apical prior" represents an architectural bias that is functionally useful for contextual integration in neural networks.
 To this end, we implemented (1) a neural network architecture with distinct apical compartments and a neuron-wise integration rule based on the observation that apical dendrites are gain modulators of somatic activity (Waters et al., 2003; Larkum et al., 2007), and developed (2) an ambiguous image classification task that requires contextual integration to be solved.
 We trained (1) the contextual modulator (top-down synapses onto apical dendrites) by gradient descent to simultaneously learn all input-context conditions defined in our task.
-We found that this bio-inspired dendritic architecture outperformed a single-compartment model which processes both sensory and contextual information in the compartment.
-In this task setting, we find there is an inherent benefit to have contextual information propagated via a distinct apical compartment.
-By applying LRP, we find that the learned top-down modulation of pyramidal neurons is consistent with the principles of biased competition (Spratling, 2008).
-Specifically, by analyzing the apical amplitude on class-specific neurons for ambiguous stimuli, we found the highest gain modulation in class-specific neurons that matched the target class.
-Moreover, we found that the level of competition introduced by the context adapted to the level of ambiguity in the task: The top-down network learned to a) refine ambiguous representations by applying strong but sparse gain modulation to basal signals when contextual information is both available and relevant, and b) provide weak modulation in the case of contradictory or irrelevant contexts or unambiguous stimuli.
+In this setting, we found that this bio-inspired dendritic architecture outperformed a single-compartment model which processes both sensory and contextual information with a single non-linear activation.
+By applying LRP, we analyzed the amplitude of apical inputs onto class-specific neurons for ambiguous stimuli.
+We found the highest gain modulation in context-relevant neurons (matching the target class), consistent with the principles of biased competition (Spratling, 2008).
+Moreover, we found that the competition introduced by the contextual integration adapted to the level of ambiguity in the task: The top-down network learned to a) refine representations of ambiguous stimuli by applying strong but sparse gain modulation to basal signals when contextual information is both available and relevant, and b) provide weak and non-specific modulation in the case of unambiguous stimuli and contradictory or irrelevant contexts, thus preserving the bottom-up representations.
 3.1 Comparison with existing computational frameworks
@@
 one-hot encoded task id) is always assumed to be relevant.
-Alternative approaches to flexible contextual integration have involved explicitly learning to prevent associations between irrelevant context and sensory inputs ((Baronig and Legenstein, 2024)), which represents a successful example of how to handle spurious contexts in a simple task.
-As an extension of this approach, our model must infer how to handle the interaction between sensory input and context across distinct tasks and scenarios, where context is relevant, irrelevant, or contradictory.
-Crucially, we only found a significant advantage of the apical prior over other integration approaches in the scenario where inputs are unambiguous and context is contradictory (and should be ignored), with an accuracy increase of 1.9% (MNIST) for models with distinct apical compartments.
-This highlights a specific advantage for apical dendrites in the role of contextual integration, where contextual relevance is crucial to performance.
-Furthermore, while the task can be solved using either the Hadamard or additive rule (Appendix 4), we found that, under uncertainty, the Hadamard integration rule enabled the model to ignore contradictory contextual signals more effectively compared to additive integration.
-Another key difference with most previous work is that as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike Wybo et al.
+Alternative approaches to flexible contextual integration have involved explicitly learning to prevent associations between irrelevant context and sensory inputs (Baronig and Legenstein, 2024), which represents a successful example of how to handle spurious contexts in a simple task.
+As an extension of this approach, our model must infer how to handle the interaction between sensory input and context across distinct tasks and scenarios, where context is relevant, irrelevant, or contradictory, but without explicit knowledge of the task.
+Crucially, we only found a significant advantage of the apical prior over single compartment integration in the scenario where inputs are unambiguous and context is contradictory (and should be ignored).
+This highlights a specific advantage for apical dendrites in the role of contextual integration, where accounting for varying contextual relevance is crucial to broadly robust performance.
+Furthermore, while we show that the task can be solved using either a gain modulation (Hadamard) or additive apical integration rule (Appendix 1 Table.
+4), we found that, under low contextual certainty, the Hadamard integration rule (apical prior) enabled the model to ignore contradictory contextual signals more effectively compared to additive apical integration.
+Another key difference of our study with most previous work is that as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike Wybo et al.
 (2022) and (Naumann et al., 2022), but inline with experimental evidence (Larkum et al., 1999).
-From an efficiency perspective, this maintains the sparsity of the network.
-One notable prior work which applies context-dependent modulation to feedforward networks is context-dependent gating (XdG) (Masse et al., 2018), which could be comparable with our basal-only ablation as it is a feedforward architecture.
-The main difference with our basal-only ablation is that XdG a) is not informed by intermediate sensory representations, and as most prior works, assumes context always agrees with sensory input.
-Therefore, its is unlikely that XdG, implemented as-is, can deal with cases where context should be overridden by sensory input, as in the unambig.Contra case.
-The advantage of apical dendrites are clear in our multi-scenario contextual integration task.
-In order to compare our results with previous work, we examined the effect of our integration mechanism in a standard multi-task setting (given task-ids).
-Under these conditions we found that that all variants perform similarly well.
-This can be explained by the setting being simpler, whereas the distinct apical compartment shines under more varied and challenging types of contextual integration that require inference of ambiguity.
-While not explicitly modular, our framework could be seen as in accordance with the Global Workspace Theory along with the "information bottleneck" (to discard irrelevant information, which can include contextual information) Goyal et al.
+Thus, our approach maintains the sparsity of the network representations, which can preserve selectivity, and would also be more metabolically efficient.
+One notable prior work which applies context-dependent modulation to feedforward networks is context-dependent gating (XdG) (Masse et al., 2018), which implements contextual inputs by concatenation of context signals and gating task-specific neurons.
+The former could be comparable with our single compartment (basal-only) ablation as it employs a single zone of integration.
+The main difference with our basal-only ablation is that XdG a) is not informed by intermediate sensory representations, and b) as most prior works, assumes task labels are given and that context always agrees with sensory input.
+Therefore, it is unknown if XdG, implemented as-is, can deal with cases where context should be overridden by sensory input, as in the unambig.Contra case.
+In our experiments, the advantage of apical dendrites was revealed in our multi-scenario contextual integration task.
+In order to compare our results with previous work, we examined the effect of our integration mechanism in a standard multi-task setting (with task-ids provided to the network, 4).
+Under these conditions we found that all variants perform similarly well.
+This can be explained by the setting being simpler, whereas the distinct apical compartment shows its advantage under more varied and challenging types of contextual integration that require inference of ambiguity (no task ids provided).
+While not explicitly modular, our framework could be incorporated into the Global Workspace Theory (see (Goyal et al., 2022) for a comparison of recent implementations), to provide a candidate mechanism for resolving inconsistency between modules under ambiguity.
+It is important to note that our architecture embodying the apical prior presupposes there is asymmetry for how contextual (top-down) and perceptual (bottom-up) representations are combined under ambiguity, which is included in some (e.g.
+Mittal et al.
+(2020)) but not other Goyal et al.
+(2022) implementations of Global Workspace Theory.
+In general, modular representations can be learned in a distributed and scalable manner, with interactions between local populations (or modules).
+As highlighted in Goyal et al.
+(2022), this frees up the capacity of a learning system for more tasks, as it can more efficiently identify task-relevant information and reuse what it has learned.
+The "apical prior" here implies that these interactions have an asymmetry: they can be contextual signals (in our case via top-down projections to apical compartments), or feed-forward/bottom-up signals.
+Here we find that this asymmetry provides flexibility to solve perceptual tasks under diverse ambiguity and contextual scenarios.
+3.2 What is necessary for robust contextual integration?
+Thus far, we have shown that in our contextual integration task, the apical prior provides a specific advantage compared to single-compartment integration when sensory or contextual signals are unreliable.
+Here, we extend these findings to make broader predictions for what specific components/computations might be necessary for robust contextual integration, and how such a mechanism could plausibly be implemented in the cortex.
+Our results suggest that it is sub-optimal to sum all activity together in a single (basal) compartment, if sensory information and context can be contradictory, as would be the case for general cognitive systems.
+This highlights a key difference between single and two-zone integration of information.
+As such, an interesting experiment would be to compare the single vs two-zone integration of bottom-up representations and top-down contextual representations in in-vivo experiments or simulations of neurons which more faithfully respect the biophysical constraints of pyramidal neurons in the neocortex, to observe whether our results hold in a more biologically plausible setting, or if it is just a feature of ANNs and/or how they were trained here.
+Furthermore, based on our results, we can infer that the apical compartment encodes some information about the ambiguity of the sensory input, and the relevance of context for a given sensory input.
+An interesting experiment could involve silencing or deactivating apical dendrites in more realistic neurons (biophysical simulations or in-vivo experiments), to assess the effect on accurate perception for our proposed contextual integration tasks.
+Our analysis showed that mutual information between apical and basal activation is positive for the ambig.Match scenario.
+On the other hand, mutual information is close to 0 for the unambig.Irrel and unambig.Contra scenario (Appendix, Table 3).
+This is not surprising given we explicitly trained under the assumption that local representations, given matching unambiguous contextual information, should collapse to the target representation.
+Moreover, when sensory ambiguity is low, local representations should be invariant to contextual ambiguity, and mismatching contextual information.
+This is an important feature of the model supporting robust contextual integration given that most available contextual signals in rich, multi-modal real-world sensory experiences are distractors for any one particular task, and should be discarded.
+Following this, we predict that a model cannot learn the relevance of context without encountering and recognizing cases with irrelevant context during training (negative samples).
+Extending this idea to unsupervised learning, it is interesting to think about how the model could identify contextual relevance.
+In an unsupervised setting this could arise from an information-maximization learning rule like InfoNCE, which uses a contrastive loss with positive and negative pairs defined via augmentation of the input (positive, relevant) against other inputs (negative, irrelevant).
+Unfortunately, the creation of positive and negative pairs relies heavily on modality-specific knowledge to construct useful data augmentations (e.g.
+a rotation applied on an image).
+A more general approach will be necessary to create such contrastive pairs for many modalities in parallel.
+An important feature of our model is that it could learn a mapping of top-down signals onto the apical compartment separately from the contextual representations.
+This is compatible with regions high in the cortical hierarchy, such as the hippocampus (via the entorhinal cortex), with independent mechanisms for learning a) how they represent contexts, and b) how those contextual representations modulate the sensory representations of lower regions (Maren et al., 2013).
+Because we assume the stability of the sensory and contextual representation prior to training contextual integration, we cannot make any conclusions from the present study about how these two representations could be simultaneously co-learned.
+However, we speculate that in the neocortex, there should be some degree of stability in the sensory and contextual representations before learning to perform contextual integration.
+This could be compatible with the distributed learning of modules (cortical regions) for sensory and contextual representations initially via self-supervised learning, prior to learning the contextual interactions between the modules (carried by cortico-cortical white matter projections).
+3.3 Limitations
+While our model unifies the function of contextual integration with key biophysical properties of dendrites, some of our modeling assumptions and simplifications are worth discussing with respect to their empirical support.
+Firstly, unlike other computational frameworks such as George et al.
+(2020) which seek to incorporate the rich diversity and laminar architecture of the neocortex, our model is designed to implement some prominent features of cortical pyramidal neurons without focusing on a specific cortical layer, such as layer 2/3 or layer 5 pyramidal neurons.
+Specifically, we focus on the functional properties of two-zone dendritic integration.
+In the cortical circuit, these mechanisms appear to be conserved across L2/3, L5, L6 pyramidal neurons (Larkum et al., 1999; Waters et al., 2003; Larkum et al., 2007; Ledergerber and Larkum, 2010).
+Here, we assume that contextual signals are provided by neurons in a higher-order region, which would likely correspond to outputs of layer 5 pyramidal neurons in that region, and integrated in the apical dendrites of putative layer 2/3, 4 or 5 pyramidal neurons (Bastos et al., 2012; Schuman et al., 2021).
+It is worth noting that top-down projections may also innervate the basal compartment of layer 2/3 neurons, and the lower regions of L5 pyramidal neuron basal dendrites (Harris et al., 2019).
+Despite these missing anatomical features, the framework developed here could serve to study the functional role of laminar architecture in the neocortex: Our study relies largely on the functional mechanism of two-zone integration in pyramidal neurons, which appears to be consistent across pyramidal neuron types, despite potential differences in the sources of the inputs they receive.
+Furthermore, our model only implements top-down modulation at a single level in the hierarchy.
+An obvious next step would be to develop a hierarchical generalization of the approach.
+While it is unknown how much extra improvement this would provide, evidence of top-down neuron masking (a less expressive form of modulation) at every layer has been shown to improve out-of-distribution robustness (Liu et al., 2023).
+One potential criticism is that we train here using gradient descent, for which concerns have been raised as to its biologically plausibility (Lillicrap et al., 2020).
+The aim of this study is not to find biological learning rules, but rather to understand contextual integration.
+We use gradient descent as a "best in class" general learning approach to find good solutions to our task, which allows us to analyze the learned representations and the respective contributions of apical and basal compartments.
+Our framework lacks the implementation of biological plasticity rules.
+We leave to future work the exploration of solutions obtained via more biologically plausible plasticity rules such as Lee et al.
+(2015); Meulemans et al.
+(2022), compared to those we obtained via backprop.
+Indeed the analysis above suggests that information-maximization plasticity rules between basal and apical compartments could be a promising direction.
+3.4 Outlook
+Often in human brains, inputs from different modalities are processed simultaneously.
+As a future direction, we propose that our model could be extended to hierarchies of moduels of sensory modalities interacting with each other.
+Deep learning models that implement multi-modal interactions are an active area of research (Radford et al., 2021; Alayrac et al., 2022; Mustafa et al., 2022).
+While they excel at multi-modal generation conditioned on one modality, they struggle with conflicting information or contradictions in mixed-modality data (Chen et al., 2024).
+As state of the art deep learning models increase exponentially in size, many recent works focus on scalable, parameter-efficient fine-tuning for target tasks, a prominent example being low-rank adapters (LoRAs) Hu et al.
+(2021); Chen et al.
 (2024).
-Under both frameworks, representations can be learned in a distributed and scalable manner, and mapped to local populations as context signals (in our case via apical compartments), and through biased competition, integrated with bottom-up representations to solve perceptual tasks.
-As highlighted in their work, this frees up the capacity of a learning system for more tasks, as it can more efficiently reuse what it has previously learned and identify task-relevant information, compared to single-compartment, single-integration models.
-3.2 Predictions: What is necessary for robust contextual integration?
-Until now, we have shown that in our contextual integration task, the apical prior provides a specific advantage compared to single-compartment integration when sensory or contextual signals are unreliable.
-Here, we aim to extend these findings to make broader predictions for what specific components/computations might be necessary for robust contextual integration, and how such a mechanism could plausibly be implemented in the cortex.
-Our results show that, given both sensory input and contextual input must be considered simultaneously, it may be sub-optimal to sum all activity together in the basal compartment, when sensory information and context are contradictory.
-At a higher level, this highlights a key difference between single and two-zone integration of information.
-As such, an interesting experiment would be to compare the single vs two-zone integration of bottomup representations and topdown contextual representations in simulations of neurons which more faithfully respect the biophysical constraints of pyramidal neurons in the neocortex, to observe whether our results hold in a more biologically plausible setting, or if it is strictly a feature of ANNs.
-Furthermore, based on our results, we can infer that the apical compartment encodes some information about the ambiguity of the sensory input, and the relevance of context for a given sensory input.
-An interesting experiment could involve silencing or deactivating apical dendrites in more realistic neurons (biophysical simulation or biological), and observing the effect on accurate perception in a similar contextual integration task.
-Mutual information between apical and basal activation is positive for the ambig.Match scenario, where we see a reduction in entropy of the prediction after observing the contextual signal.
-On the other hand, mutual information is near 0 for the unambig.Contra scenario, which aligns with our main findings (see Appendix, Table 3).
-This is not surprising given that we performed supervised training.
-We explicitly trained under the assumption that local representations, given mis-matching context, should be invariant to sensory ambiguity.
-This could be interpreted as being surprised by a stimulus, but not having informative or relevant context to reduce the ambiguity of the stimulus.
-This is important for robust contextual integration if one subscribes to the hypothesis that, most available contextual signals in rich, real-world sensory experience are distractors for any one particular task, and should be discarded.
-Following this, we predict that a model cannot learn the relevance of context with out encountering scenarios with irrelevant context during training.
-As an ablation study, we could remove the unambig.Irrel scenario from our loss function.
-We expect to observe a significant drop in accuracy on that scenario, reflecting the inability to assess contextual relevance.
-Extending this idea to unsupervised learning, it is interesting to think about how the model could identify contextual relevance without a supervised signal and be invariant to sensory ambiguity under unhelpful context.
-In an unsupervised setting this could arise from an information-maximization learning rule like InfoNCE, which uses a contrastive loss with positive and negative pairs defined via augmentation of the input (positive) against other inputs (negative).
-Unfortunately, the creation of positive and negative pairs relies heavily on domain-specific knowledge for useful data augmentations (eg.
-a rotation applied on an image).
-If it is possible to create contrastive pairs in a domain-agnostic manner, such an unsupervised learning framework could approximate the effects of supervised training shown in our results.
-Given the fact that most associations of context and input are either irrelevant or contradictory, and only a small subset of associations represent matching context, it is not surprising that the gain modulation is sparse.
-We expect also that our high-dimensional apical weight matrices could be well approximated with low-rank matrices without losing significant performance, in line with work in modulating deep learning models for downstream tasks via low-rank adapters Hu et al.
-(2021).
-While this method uses resources efficiently, it lacks the mechanism to separate irrelevant or contradictory from useful contexts that the apical prior excels at, which indicates a potential synergy between the two.
-We hypothesize the brain uses both mechanisms to learn to handle the large pool of available sources of context efficiently.
-A future direction could combine LORAs with the apical prior for contextual integration tasks.
-We find that our model could learn a mapping of top-down signals onto the apical compartment, which importantly, was learned separately from the contextual representations.
-This is compatible with high-order cortical regions such as the hippocampus, where the learning and updating of contextual representations is disentangled from how they modulate sensory representations (Maren et al., 2013).
-Because we assume the stability of the sensory and contextual representation, we cannot make any conclusions about how these two representations could be simultaneously co-learned.
-However, we speculate there should be some degree of stability in the sensory representations to learn to perform contextual integration.
-3.3 Limitations
-While our model unifies the function of contextual integration with key biophysical properties of dendrites, some key model assumptions and simplifications are worth discussing with respect to their empirical support.
-Firstly, our model is designed to implement some key features of cortical pyramidal neurons without focusing on a specific subpopulation that make up cortical layers, including layer 2/3 and layer 5 pyramidal neurons.
-Here, we assume that contextual signals are provided by neurons in a higher-order region, which would likely correspond to layer 5 output, and integrated in apical dendrites of putative layer 2/3 pyramidal neurons (Bastos et al., 2012; Schuman et al., 2021).
-It is noteworthy though that layer 5 neurons can also send top-down projections to the basal compartment of layer 2/3 neurons, although to a lesser extent (Harris et al., 2019).
-Unlike other computational frameworks such as (George et al., 2020) which seek to incorporate the rich architectural diversity of the neocortex, we sought to simplify and abstract many features (using neural networks) in order to be able to train models to solve more challenging contextual integration tasks which could give useful insights that could not be seen with simpler tasks.
-Our model heavily abstracts temporal dynamics that are normally observed in biological circuits.
-For example, in the presence of ambiguous stimuli, human subjects can oscillate between two interpretations, a phenomenon termed ’perceptual bistability’ (Hardstone et al., 2021).
-During perceptual bistability, neural activity alternates between discrete states (Abeles et al., 1995; Akrami et al., 2009), following principles of continuous attractor dynamics (Daelli and Treves, 2010).
-3.4 Outlook
-In the future, our model could support more realistic temporal dynamics, which would enable the testing of a rich array of hypotheses regarding contextual integration.
-In the future, we propose that our model could be extended to hierarchies of sensory modalities interacting with each other.
-Deep learning models that implement multi-modal interactions are an active area of research (Radford et al., 2021; Alayrac et al., 2022; Mustafa et al., 2022).
-While they are very good at multimodal generation with conditioning on one modality, they do not currently have the ability to resolve conflicting information and can be easily confused by contradictions in mixed-modality data.
-While these models employ transformer architectures that integrate context differently than MLPs, our loss function could inspire new approaches for robust multi-modal perception in deep neural networks.
-From a neurophysiology perspective, our framework could be leveraged to test specific hypotheses of contextual integration in health and disease, and in particular in cases where the physiological properties of pyramidal neurons are known to be altered (Kim et al., 2014).
+While LoRAs use resources efficiently, they lack a contextual integration component, and as such, they could be complementary to the "apical prior" architecture studied here.
+A future direction could combine LoRAs with the apical prior for efficient contextual integration tasks at scale.
+For example, given that in our task, only a small subset of associations represent matching context, we expect that the weight matrices mapping contextual representations onto apical dendrites could be effectively replaced by LoRAs.
+This has better scaling properties due to using a smaller fraction of total network capacity.
+Initially, stable and useful representations of context could be achieved with self-supervised learning.
+We speculate that the brain might employ both the apical prior and sparse weights to learn to handle the large pool of available sources of context efficiently.
+Toward Global Workspace Theory, we argue that a) there must be a mechanism for resolving ambiguity or discrepancy between modules before the step in which modules compete to write to the shared global workspace (as in Goyal et al.
+(2022)), and b) implementation of the "apical prior" satisfies this requirement.
+Thus, our present study provides an important step in understanding the process by which brain-wide networks of many modalities can coordinate and integrate information.
 4 Methods
@@
 4.2 Contextual integration model architecture
-The contextual integration model consists of three components: the pretrained backbone that computes representations of input images, the readout network for classifying their respective representations, and the topdown network which modulates the intermediate representations from the pretrained backbone based on contextual information.
-The backbone is implemented as a convolutional variational autoencoder (VAE), and was pretrained on unambiguous characters with the standard ELBO loss.
+The contextual integration model consists of three components: the pre-trained backbone that computes representations of input images, the readout network for classifying their respective representations, and the top-down network which modulates the intermediate representations from the pre-trained backbone based on contextual information.
+The backbone is implemented as a convolutional variational autoencoder (VAE), and was pre-trained on unambiguous characters with the standard ELBO loss.
 VAEs are known to learn a smooth latent space, which we assume here to be an important ingredient to allow learning of top-down contextual modulations.
 Next, we freeze the VAE weights and train a readout (MLP) on the VAEs latent representations of images with a classification objective (cross entropy loss).
-Finally, we freeze these components and train the topdown network in two contextual tasks, with oracle and temporal context respectively, as described in the next sections.
-The topdown network $g(\cdot)$ is implemented as a multilayer perceptron (MLP) mapping the concatenated VAE latents and context vector to apical activations (eq.
+Finally, we freeze these components and train the top-down network in two contextual tasks, with oracle and temporal context respectively, as described in the next sections.
+The top-down network $g(\cdot)$ is implemented as a multi-layer perceptron (MLP) mapping the concatenated VAE latents and context vector to apical activations (eq.
 3), which then modulate the VAEs intermediate representation.
@@
 We trained our GRU to predict the modulo 10 sum of MNIST digits in a sequence of two digits’, reflecting the task in our sequential MNIST/EMNIST dataset.
-An important detail is that the input to the GRU at time $t$ is not the image $x_{t}$ , but rather the latent $\mathbf{Ub}_{t}$, encoded by the pretrained backbone.
+An important detail is that the input to the GRU at time $t$ is not the image $x_{t}$ , but rather the latent $\mathbf{Ub}_{t}$, encoded by the pre-trained backbone.
 After processing a sequence of digits ($x_{t-2},x_{t-1}$), the output state $o_{t}$ of the GRU was then used as a top-down context signal, in place of the one-hot context from the oracle case.
@@
 4.3.1 Loss function
-The topdown network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function jointly across the following 5 conditions: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra).
+The top-down network $g(x,c|\Theta_{a})$ is explicitly trained to minimize a loss function jointly across the following 5 conditions: ambiguous input with matching context (ambig.Match), unambiguous input with matching context (unambig.Match), unambiguous input with irrelevant context (unambig.Irrel), ambiguous input with irrelevant context (ambig.Irrel), and unambiguous input with contradictory context (unambig.Contra).
 We use mean squared error as the loss criterion between the predicted (computed as in equation 2) and the target latent representations.
-Each minibatch is sampled from dataset $D$, and training data for each condition are equally balanced in the minibatch.
+Each mini-batch is sampled from dataset $D$, and training data for each condition are equally balanced in the mini-batch.
 $\mathbf{L}(\Theta_{a};D)=\mathbbm{E}_{((x_{0},y_{0}),x_{ambig},(x_{1},y_{1}))\sim D,c_{match},c_{irrel}\sim C}[\sum_{s=1}^{5}||\mu_{s}^{*}-\hat{\mu_{s}}||_{2}^{2}]$ (5)
@@
 […]
-3:for triplet $\texttt{triplet minibatch}(\mathbf{x_{0}},\mathbf{y_{0}}),\mathbf{x_{ambig}},(\mathbf{x_{1}},\mathbf{y_{1}})\texttt{ sampled from }D$ do
+[…]
 4:     $y\leftarrow cat(\mathbf{y_{0}},\mathbf{y_{1}})$